### v1

# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

Abstract.
Cued Speech (CS) enhances lipreading through hand coding, providing precise speech perception support for the hearing-impaired. CS Video-to-Speech generation (CSV2S) task aims to convert the CS visual expressions (CS videos) of hearing-impaired individuals into comprehensible speech signals. Direct generation of speech from CS video (called single CSV2S) yields poor performance due to insufficient CS data. Current research mostly focuses on CS Recognition (CSR), which convert video content into linguistic text. Based on this, one straightforward way of CSV2S is to combine CSR with a Text-to-Speech system. This combined architecture relies on text as an intermediate medium for stepwise cross-modal alignment, which may lead to error propagation and temporal misalignment between speech and video dynamics. To address these challenges, we propose a novel approach that directly generates speech from CS videos without relying on intermediate text. Building upon this, we propose UniCUE, the first unified framework for CSV2S, whose core innovation lies in the integration of the CSR task that provides fine-grained visual-semantic information to facilitate speech generation from CS videos. More precisely, (1) a novel fine-grained semantic alignment pool to ensure precise mapping between visual features and speech contents; (2) a VisioPhonetic adapter to bridge cross-task representations, ensuring seamless compatibility between two distinct tasks (i.e., CSV2S and CSR); (3) a pose-aware visual processor is introduced to enhance fine-grained spatiotemporal correlations between lip and hand movements in CS video. Experiments on our new established Chinese CS dataset (14 cuers111Cuer represents the person who perform CS.: 8 hearing-impaired and 6 normal-hearing) show that our UniCUE significantly reduces Word Error Rate by 78.3% and improves lip-speech synchronization by 32% compared to the single CSV2S.
1. Introduction
Cued Speech (CS) is an advanced visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010). CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (Cornett, 1967; Leybaert and LaSasso, 2010; Leybaert et al., 2010). In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels, addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019), as illustrated in Figure 1(a).
CS Video-to-Speech generation (CSV2S) task aims to convert the visual expressions (CS videos) of hearing-impaired individuals into comprehensible speech signals. However, directly constructing an end-to-end CSV2S model faces the following challenges: First, the complex semantic correlation between CS video and semantic content requires precise mapping from visual cues (lip reading and hand coding) to speech, while limited CS dataset scale further constrains model capacity. Second, fine-grained visual spatiotemporal modeling is crucial to resolve the natural asynchrony where hand gestures initiate earlier than corresponding lip movements i.e., hand-preceding phenomenon (Liu et al., 2020). Third, personalized adaptation must address dual variations: (1) inter-cuer differences in CS expression, and (2) acoustic diversity in cuer identity characteristics. To the best of our knowledge, no prior work has addressed CSV2S task.
Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text(Liu et al., 2020, 2024, 2018, 2019), neglecting the critical need for natural speech generation. This gap severely limits real-time communication between hearing-impaired and normal-hearing individuals, particularly in educational and social contexts. For example, in social gatherings, normal-hearing individuals may need to quickly and accurately understand and respond to questions or information shared by their hearing-impaired friends. However, existing CSR methods can only provide text output, which is insufficient to support natural and smooth interactions. Moreover, lip-reading video-to-speech synthesis methods like LipVoicer (Yemini et al., 2023) leverage lip motions alone, and they fail to model the complementary hand coding in CS. Thus, the primary motivation of this study is to build the first CSV2S system that directly decoding Chinese CS video into speech, as shown in Figure 1(b).
Based on the existing CSR methods, a straightforward approach to build CSV2S is to combine CSR with a Text-to-Speech (TTS) model, as shown in Figure 2(a). However, this combined architecture has fundamental drawbacks. Firstly, the intermediate CS text serve as a medium between visual and speech, inheriting recognition errors from the CSR module, leading to inaccurate speech content. Secondly, the CSR stage only extracts semantic text but loses fine-grained temporal information from video, resulting in generated speech lacking temporal alignment.
Recent advances in multimodal generation have demonstrated that integrating semantic reasoning capabilities from vision-language models (VLMs) can significantly enhance complex generation tasks, such as text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a). Inspired by this paradigm, we believe that the fine-grained visual understanding inherent in CSR can serve as a semantic compensator for CSV2S generation. As illustrated in Figure 2(b), our framework introduces a shared visual processor that bridges CSR and CSV2S tasks. This processor operates as a two-way translator: during CSR, it extracts linguistic semantics from spatiotemporal lip-hand patterns; for CSV2S, it maps these semantics to guide speech synthesis while preserving cuer identity. The key innovation lies in explicitly modeling the semantic compensation flow, where CSR-derived phoneme-level constraints reduce potential ambiguities in word-level speech generation, overcoming the challenges of fine-grained semantic correlation modeling.
To address the above challenges, building upon the semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, Unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently, relying solely on raw video embeddings, UniCUE is equipped with a pose-aware visual processor that simultaneously processes both modalities. This allows UniCUE to generate a mixed visual representation by combining video embeddings and pose embeddings. This mixed setup not only achieves fine-grained spatiotemporal modeling for hand-preceding phenomenon but also enhances robustness to cuer-specific CS expression. Secondly, to enhance the semantic alignment between visual cues and spoken content, we introduce a semantic alignment pool. This mechanism aligns the video latent space and the pose latent space with linguistic semantics via contrastive learning, enabling the model to capture complex semantic correlations and improve the accuracy of the generated speech. Thirdly, we incorporate the visual processor from CSR into the diffusion-based CSV2S model, using an VisioPhonetic Adapter (VPA) to translate fine-grained representation into diffusion-compatible codes. Based on the identity correlation between speech and face portrait (Rong and Liu, 2024), this design enables our diffusion decoder to not only effectively integrate the semantic information conveyed by the visual cues but also capture the unique characteristics of the cuer identity in the visual frames, thereby enhancing accurate cuer-specific speech generation.
To evaluate UniCUE on hearing-impaired individuals, we collect CS data from 8 hearing-impaired cuers. Additionally, we extend the Mandarin Chinese CS dataset (Lei et al., 2024b) by incorporating data from 8 hearing-impaired and 6 normal-hearing cuers (individuals who perform CS), resulting in a total of 11,282 CS videos. The results on this dataset demonstrate that our method not only achieves accurate linguistic understanding but also preserves the cuer’s characteristics.
The main contributions of this work can be summarized as:
We propose the first CSV2S generation framework by constructing a new unified multimodal system that integrating CSR capabilities to enhance CSV2S.
We propose a pose-aware visual processor and a semantic alignment pool to enhance fine-grained, semantically aligned visual representations, and introduce an VPA module to convert fine-grained semantic information into understandable coding for the speech synthesis model.
We establish a new Mandarin Chinese CS dataset comprising 8 hearing-impaired and 6 normal-hearing cuers. Experimental results demonstrate that our UniCUE framework outperforms the single CSV2S model in terms of speech consistency, accuracy, and time efficiency.
2. Related Work
2.1. Video-to-Speech Generation
V2S involves generating natural-sounding speech synchronized with a silent video of a person talking. A key challenge in this area is performing complex visual-speech modeling under the limited data. To address this challenge, Uni-Dubbing (Lei et al., 2024a) incorporates modality alignment during the pre-training phase on multimodal datasets and then fine-tunes using both multimodal and single-modality audio data. Similarly, Kefalas et al. (Kefalas et al., 2024) leverage abundant audio-only datasets to pre-train the speech encoder and decoder, subsequently fine-tuning with paired speech-video data. Instead of relying on external data, some works (Kim et al., 2023; Yemini et al., 2023; Gupta et al., 2024) have found that incorporating text transcripts plays a significant role in enhancing V2S generation. For example, Kim et al. (Kim et al., 2023) propose guiding the model using multimodal supervision, i.e., text and speech, to complement insufficient word representations via multi-task learning. In contrast to directly using text transcripts, we innovatively utilize the shared semantic correlations between visual-text and visual-audio in CSR and CSV2S.
2.2. Cued Speech Recognition
CS is an augmented lip-reading system enhanced by hand coding, offering significant support to the hearing-impaired community. While CSR task focuses on converting CS videos into linguistic texts, leveraging both lips and hands as distinct modalities that provide complementary information (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021). Existing CSR methods often process lip and hand features independently, followed by feature fusion for content recognition (Liu et al., 2020, 2024; Liu and Liu, 2023; Zhang et al., 2023). Due to the asynchronous nature of lip and hand movements, achieving effective fusion remains a challenging problem. Liu et al. (Liu et al., 2020) introduced a re-synchronization procedure to align hand features with lip features, addressing the temporal mismatch. While a transformer-based cross-modal mutual learning framework has been developed to facilitate multimodal interaction (Liu and Liu, 2023; Liu et al., 2024). Zhang et al. (Zhang et al., 2023) further explored privacy risks in the training process by applying federated learning. In contrast, our approach takes whole video frames as input that models lip and hand features simultaneously, avoiding the complexities of feature fusion. Moreover, we propose a pose-aware visual processor to enhance feature representation between the two modalities, thereby improving recognition accuracy.
2.3. Unified Understanding and Generation
Recent efforts in unifying understanding and generation tasks can be broadly categorized into two paradigms. The first paradigm integrates visual-language understanding with external generative models (e.g., diffusion models) to achieve multimodal generation (Wu et al., 2023; Dong et al., 2023; Jin et al., 2023; Li et al., 2024; Sun et al., 2024; Ge et al., 2024b, a). For instance, works like (Jin et al., 2023; Li et al., 2024) leverage large language models (LLMs) for semantic comprehension while relying on diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-quality image synthesis. The second paradigm focuses on optimizing a LLMs-based foundation models through next-token prediction objectives for both visual understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2024a; Chen et al., 2025b). For instance, Transfusion (Zhou et al., 2024) unifies the processes of image understanding and image generation within a transformer. This approach not only inherits the strong comprehension ability of language models but also preserves the detailed information of images, achieving better controllable generation results in text-to-image. Despite their success, these frameworks predominantly address visual-text scenarios, leveraging understanding for speech generation is unexplored. In contrast to prior works, we present the first unified framework that bridges visual understanding and speech generation.
3. Method
3.1. UniCUE
To achieve accurate CSV2S generation, the proposed framework need simultaneously resolve two critical challenges: (1) semantic understanding of the semantic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment. Inspired by the auxiliary benefits of unified understanding and generation for multimodal controllable image synthesis (Wu et al., 2023; Dong et al., 2023), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations. As illustrated in Figure 3, the framework operates via two pathways.
CSR: Fine-grained Visual Cues Understanding. Given a CS video and its corresponding pose (extracted via OpenPose (Cao et al., 2019)), a pose-aware visual processor extracts the embedding of the visual mixture . The visual embedding is first transcribed into linguistic text via a transformer-based text decoder to model a fine-grained spatiotemporal visual representation space. The process can be denoted as: , where is the predicted text tokens. The learning objective of the CSR path is to predict text tokens as closely as possible to the ground truth text tokens :
where and are the -th token of the predicted tokens and ground truth, respectively, and is the total number of tokens. By transcribing cross-cuer visual cues into linguistic text, CSR path establishes phoneme-level semantic grounding for the visual embedding , which equips with cuer-adaptive, fine-grained semantic understanding capabilities.
CSV2S: Cuer-specific Speech Synthesis. Based on the correlation between face portrait and speech (Rong and Liu, 2024), a latent diffusion model (LDM) (Rombach et al., 2022) is trained to synthesize cuer-specific speech conditioned on the learned visual embedding . Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes , which are then corrupted into by sequentially injecting Gaussian noise over time steps:
where represents the noise level at the time step. The noised is then input into the LDM (Rombach et al., 2022) to learn the distribution of latent speech codes under the visual condition . The objective of the CSV2S generation path can be formulated as:
where represents the LDM. By training LDM, visual information is interacted with latent speech codes, enabling the generated speech to accurately match the dynamic changes in the visual content while preserving the naturalness of personalized speech.
UniCUE: Unified Understanding and Generation. The CSR path learns fine-grained visual embeddings by performing linguistic text recognition, which are then used as conditional inputs for the CSV2S path, enabling CSV2S to leverage the fine-grained visual understanding for accurate speech generation. This shared mechanism in our unified framework not only reduces information loss but also avoids error propagation that might arise from intermediate text conversions. Therefore, CSV2S can capture cue-specific acoustic features while maintaining linguistic accuracy and temporal consistency, generating cuer-specific speech outputs.
3.2. Pose-aware Visual Processor
Considering the spatiotemporal correlation between hand coding, lip reading, and their semantic content, both CSV2S and CSR require precise localization of lip and hand keypoints. This necessity motivates the visual encoder to capture fine-grained pose features. To address this, we develop a pose-aware visual processor to construct mixed visual representations by synergistically integrating pose and video embeddings, as shown in Figure 3.
Specifically, given the video input and pose input , a tokenizer first splits them into patch tokens. Let the ViT-based image encoder be denoted as . We extract the video features and pose features from the given video tokens and pose tokens , respectively:
where and . Here, represents the number of visual patch tokens and denotes the dimension of visual embedding. The video features and pose features are then concatenated along channel dimension, apply a trainable projector layer to form the mixed visual embedding :
where MLP refers to the projection layer. The CSR path supervises the training of the pose-aware visual processor, guiding it to adapt to cuer-specific CS expressions and learn fine-grained semantic representations.
3.3. Semantic Alignment Pool
To further enhance fine-grained semantic consistency between visual representation and linguistic content, we propose a semantic alignment mechanism that synchronizes visual, pose, and textual modalities through contrastive learning. Specifically, a ViT-based text encoder first projects tokens of ground-truth transcripts into embeddings , sharing the same latent space with visual and pose features extracted by the pose-aware processor. We formulate the alignment objective as a multi-modal embedding alignment Loss:
where Cos(,) denotes the cosine similarity. The total semantic alignment loss is calculated via:
By individually aligning video and pose features with text embeddings, the model can effectively integrate complementary cues, enhancing its ability to understand visual cues for speech generation.
3.4. VisioPhonetic Adapter
The VisioPhonetic adapter (VPA) acts as a critical translator between the learned pose-aware visual processor in CSR and the diffusion-based speech generator in CSV2S, addressing the representation gap caused by their heterogeneous architectures. As illustrated in Figure 4, this lightweight module employs a sequential architecture to progressively refine visual-semantic representations into diffusion-compatible conditioning signal:
which includes two MLPs and a Q-Former-style (Li et al., 2023) cross-attention layer. The cross-attention mechanism is specifically designed to bridge the representation gap through learnable semantic queries. Formally, given the mixed visual embedding , the process is formulated as: where denotes the learnable query vector initialized from the diffusion model’s latent space, and , , are projection matrices.
The VPA effectively aligns visual information with the semantic requirements of speech generation, further mapping the semantic connections between hand-lip dynamics, phonetic constraints, and speech. temporally related information from the visual features, thereby improving the accuracy, synchronization, and cuer-specific expression in speech.
3.5. Training and Inference
The training process is divided into two sequential phases to ensure effective knowledge transfer from visual understanding to speech synthesis. Stage 1: CSR focuses on learning fine-grained spatiotemporal representations by jointly optimizing the pose-aware visual processor, the semantic alignment pool, text encoder and decoder. This stage trains the model using the cross-entropy loss (as defined in Eq. 1) and the contrastive alignment loss (as defined in Eq. 7), formulated as: Stage 2: CSV2S leverages the learned visual representations from Stage 1 to train the VPA network and diffusion model with the objective (denoted in Eq. 3).
During inference, the framework operates in a streamlined single-path mode to synthesize speech directly from cued CS videos. As illustrated in Figure 5 (b), given an CS video and its corresponding pose, the trained pose-aware visual processor extracts the visual mixture representation , which encodes both semantic content and cuer characteristics. A noise sample , sampled from a standard Gaussian distribution , is fed into the trained LDM to predict the speech latent code conditioned on the visual mixture representation over timesteps:
Subsequently, the predicted speech latent code is decoded into a speech mel-spectrogram via VAE decoder, and a pre-trained Vocoder (Liu et al., 2023) is used to convert the mel-spectrogram into a speech waveform. Through this series of steps, the CSV2S system is able to generate natural speech that is consistent with the cuer’s characteristics based on the input CS video and its corresponding pose information.
| Dataset | Cuers | Sentences | Resolution | FPS |
| French CS (Liu et al., 2018) | 1 normal-hearing cuer | 238 | 50 | |
| British CS (Liu et al., 2019) | 1 normal-hearing cuer | 97 | 25 | |
| Chinese CS (Lei et al., 2024b) | 4 normal-hearing cuers | 4000 | 30 | |
| Ours (Chinese CS) | 6 normal-hearing and 8 hearing-impaired cuers | 11282 | 30 | |
| Method | Normal-hearing cuers | Hearing-impaired cuers | |||||
| WER | LSE-D | LSE-C | FAD | WER | LSE-D | LSE-C | |
| GT | - | 7.250 | 0.200 | - | - | - | - |
| CMML (Liu and Liu, 2023)+TTS | 0.614 | 9.327 | 0.124 | 4.908 | 1.102 | 9.825 | 0.090 |
| CNN+FLASH (Hua et al., 2022)+TTS | 0.956 | 9.773 | 0.090 | 5.747 | 0.998 | 9.943 | 0.070 |
| CNN+Linformer (Wang et al., 2020)+TTS | 0.991 | 9.874 | 0.082 | 5.944 | 1.092 | 10.014 | 0.067 |
| CNN+Performer (Choromanski et al., 2020)+TTS | 0.945 | 9.873 | 0.105 | 5.429 | 0.988 | 9.876 | 0.079 |
| CNN+Cosformer (Zhen et al., 2022)+TTS | 1.069 | 10.063 | 0.076 | 6.355 | 1.229 | 10.115 | 0.064 |
| EcoCued (Liu et al., 2024)+TTS | 0.672 | 9.495 | 0.114 | 5.326 | 1.621 | 9.846 | 0.087 |
| Ours (CSR)+TTS | 0.030 | 8.461 | 0.158 | 3.653 | 0.254 | 9.045 | 0.127 |
| Ours (UniCUE) | 0.056 | 7.432 | 0.194 | 1.749 | 0.293 | 8.716 | 0.153 |
4. Experiment
4.1. Dataset
To further enhance the model’s generalization ability and validate its practical effectiveness in facilitating communication for hearing-impaired individuals, we establish a new Chinese CS dataset by incorporating CS data from 2 normal-hearing and 8 hearing-impaired cuers with 4 normal-hearing cuers data in (Lei et al., 2024b), increasing the total number of sentences in the dataset to 11,282. A comparison between our dataset and existing CS datasets is presented in Table 1. Specifically, we invited 2 normal-hearing cuers and 8 hearing-impaired cuers to simultaneously utter and encode a set of text sentences provided by (Lei et al., 2024b). The 10 cuers underwent systematic training to ensure they could perform Mandarin Chinese CS fluently and accurately. It is important to note that our dataset was collected with the explicit consent of all cuers and is available for open-source use. We use CS data from 6 normal-hearing cuers to train the model, splitting this data into training and test sets in a ratio to ensure effective model training and preliminary validation. To more comprehensively evaluate the practical effectiveness in facilitating communication for hearing-impaired individuals, we included all the CS data from 8 hearing-impaired cuers in the test set.
4.2. Implementation Details
Architecture Details. To construct the CSV2S path, we employ a speech generation pipeline consisting of a VAE encoder and decoder, an LDM, and a Vocoder from AudioLDM (Liu et al., 2023). We initialize the pipeline by loading the pre-trained weights of the VAE and LDM. For the CSR path, we utilize a conditional translation LLM, MBart (Liu, 2020), which comprises a tokenizer, an image encoder, a text encoder, and a text decoder. The pipeline is initialized by loading the pre-trained weights with multi-language corpus222https://huggingface.co/facebook/mbart-large-50.
Training Details. We extract CS pose information—including hands, face, and body—from videos using OpenPose333https://github.com/CMU-Perceptual-Computing-Lab/openpose. Raw video frames and pose images are resized to , while the audio is resampled to 16 kHz and converted into mel-spectrograms. The entire learning framework is implemented in PyTorch, and all experiments are conducted on an NVIDIA RTX A6000. We use the Adam optimizer with a learning rate of for end-to-end training, setting the mini-batch size to 2.
Comparison Methods. To validate the effectiveness of our framework, we conduct extensive comparisons between our method UniCUE (CSV2S with CSR-transferred information) and the following methods: 1) our single CSV2S: an end-to-end trained CSV2S without the shared CSR information; 2) our CSR: comprising of the pose-aware visual processor and text decoder to output CS linguistic text (shown in Figure 3); 3) SOTA CSR methods: including CMML (Liu and Liu, 2023), EcoCued (Liu et al., 2024), and various combinations of CNNs and transformers, such as FLASH (Hua et al., 2022), Linformer (Wang et al., 2020), Performer (Choromanski et al., 2020), and Cosformer (Zhen et al., 2022). For fairness, we reproduce these CSR methods on our dataset using the same experimental setup described in their respective papers. Since CSR methods only output phoneme-level transcriptions, we propose a two-stage conversion pipeline to synthesis speech, as illustrated in Figure 5 (a). The process consists of: 1) converting CSR outputs into Chinese pinyin sequences using rule-based CS patterns, followed by translating the pinyin into Chinese characters with QWen toolkit444https://github.com/QwenLM/Qwen;2) generating the final waveforms using SOTA TTS model F5-TTS (Chen et al., 2024) with reference speech.
| Method | Normal-hearing cuers | Hearing-impaired cuers | |||||
| WER | LSE-D | LSE-C | FAD | WER | LSE-D | LSE-C | |
| GT | - | 7.250 | 0.200 | - | - | - | - |
| CSR††+TTS | 0.125 | 8.847 | 0.129 | 4.231 | 0.307 | 9.672 | 0.098 |
| CSR‡+TTS | 0.097 | 8.716 | 0.135 | 4.063 | 0.291 | 9.472 | 0.106 |
| CSR+TTS | 0.030 | 8.461 | 0.158 | 3.653 | 0.254 | 9.045 | 0.127 |
| CSV2S† | 0.288 | 8.379 | 0.141 | 3.205 | 0.403 | 8.984 | 9.128 |
| CSV2S | 0.258 | 8.106 | 0.147 | 2.864 | 0.379 | 8.912 | 0.130 |
| UniCUE* | 0.104 | 7.741 | 9.172 | 2.219 | 0.332 | 8.884 | 0.139 |
| UniCUE | 0.056 | 7.432 | 0.194 | 1.749 | 0.293 | 8.716 | 0.153 |
4.3. Evaluation Metrics
To evaluate the performance of the synthesized speech, we employ three categories of metrics focusing on linguistic accuracy, audio-visual synchronization, and speech quality. For linguistic accuracy, we measure the Word Error Rate (WER) between the recognized text of generated speech and ground truth transcripts. The transcripts are normalized by removing punctuation and converting numbers to words prior to alignment. To measure the audio-visual synchronization between the generated speech and the corresponding video, we utilize SyncNet (Chung and Zisserman, 2016) and report two key metrics: LSE-D (the temporal distance between audio and video) and LSE-C (the confidence score). These values are computed using the pre-trained SyncNet model555https://github.com/joonson/syncnet-python. For speech quality, Fréchet Audio Distance (FAD) (Kilgour et al., 2018) is employed to assess the naturalness of generated speech by measuring the distribution divergence between real and synthesized audio in a pretrained VGGish embedding space (Hershey et al., 2017).
4.4. Comparison with SOTA Methods
We conduct comparison with SOTA methods. As shown in Table 2, our framework demonstrates superior performance: 1) Thanks to the pose-aware visual processor and semantic alignment pool, our CSR model achieves WER scores of 0.030 for normal-hearing cuers and 0.254 for hearing-impaired cuers, significantly outperforming previous SOTA CSR methods; 2) Leveraging the strong understanding capabilities of CSR, our UniCUE model also achieves lower WER scores. Additionally, our UniCUE surpasses others in FAD, LSE-D, and LSE-C metrics, showcasing high-quality, temporally aligned, and expressive speech generation.
4.5. Ablation Studies
To systematically validate the contribution of each component, we conduct controlled ablation studies on both normal-hearing and hearing-impaired test data. The results are listed in Table 3.
The key findings are analyzed as follows: 1) Unified Training Paradigm. Compared with the single CSV2S baseline, our UniCUE reduces WER by 62% (0.056 vs. 0.258) for normal-hearing cuers and 22.7% (0.293 vs. 0.379) for hearing-impaired groups. This verifies that transferring fine-grained semantic knowledge from CSR to CSV2S through shared visual representation learning effectively alleviates the challenge of complex semantic correlation.
2) Visual Processor Design. We analyze the impact of pose-aware visual processor by comparing CSR‡+TTS vs. CSR+TTS and CSV2S† vs. CSV2S. Models relying solely on raw video features exhibit suboptimal performance due to their limited capacity to capture fine-grained motion patterns. This suggests that unprocessed video frames contain substantial interference from redundant visual information. By integrating pose estimation features as supplementary inputs, the model gains enhanced fine-grained awareness of motions, leading to significant improvements in both CSR and CSV2S tasks. Furthermore, the pose features enhances model robustness against inter-cuer variability. As illustrated in Figure 6, our model maintains consistent performance across six normal-hearing cuers when speaking same linguistic content, demonstrating the importance of introducing pose-aware information. 3) Semantic Alignment Mechanism. Disabling the semantic alignment pool (CSR††+TTS) weakens the pose-text correlation, causing 17.9% relative WER increase (0.030 vs. 0.125). The contrastive alignment loss enforces spatiotemporal consistency between visual cues and CS phonemes, which is particularly beneficial for accurate semantic content. 4) VPA Mechanism. Removing the VPA (UniCUE*) results in an 85.7% degradation in FAD (1.749 vs. 2.219), indicating that the cross-attention-based VPA effectively bridges the representation gap by adaptively selecting critical visual-semantic cues relevant to speech characteristics through learnable query interaction. 5) Qualitative Analysis. To more comprehensively demonstrate the advantages of our unified framework UniCUE, we also provide qualitative comparison of CSR+TTS, CSV2S, and UniCUE. As shown in Figure 7, UniCUE achieves more accurate temporal alignment with the ground truth, indicating its superior ability to capture temporal dynamics within the visual modality. In contrast, CSR and CSV2S exhibit noticeable temporal drift.
4.6. User Study
To holistically evaluate the perceptual quality of synthesized speech, we conduct a user study. In this experiment, test samples were randomly selected from the outputs of each cuer, and 10 volunteers were invited to rate the generated speech in three perceptual dimensions on 5-point Likert scales: Accuracy (1: Unintelligible, 5: Perfect), Quality (1: Artificial, 5: Human-like), and Synchronization (1: Desynchronized, 5: Perfectly aligned). Four systems were compared: the SOTA CSR baseline (CMML (Liu and Liu, 2023)), our CSR, CSV2S, and the proposed UniCUE. Based on the mean and variance values presented in Figure 8, UniCUE demonstrates statistically significant superiority across all metrics. These results confirm that the unified framework successfully bridges the gap between visual understanding and speech generation, outperforming both combined pipelines and task-specific baselines in human perception.
4.7. Discussion and Analysis
Analysis on Cuer Representation. To quantitatively assess the capability in preserving cuer-specific acoustic characteristics, we analyze the latent cuer embeddings extracted from generated speech using Resemblyzer666https://github.com/resemble-ai/Resemblyzer. The embeddings are projected into 2D space via t-SNE (Maaten and Hinton, 2008) for visual interpretation. As shown in Figure 9, speech samples from the same cuer (e.g., cuer 1 to cuer 6) form tight cluster in the high-dimensional embedding space, while the distances between the cluster centers of different cuers are significantly larger. This indicates that the model effectively captures and distinguishes the personalized acoustic features of each cuer.
Computational Efficiency. To quantify the efficiency gains of our unified framework, we conduct time cost evaluations across both training and inference stage. (a) Training Efficiency. We compared the training performance of CSV2S and UniCUE to evaluate the training efficiency of our proposed unified framework. As shown in the Figure 10, the training loss of UniCUE decreases rapidly to a stable value in fewer epochs, indicating that the introduction of fine-grained understanding ability accelerates the learning process of the speech generation model. Moreover, the validation loss also decreases concurrently with less fluctuation, demonstrating improved model robustness. (b) Real-time Inference. To validate the real-time performance of the model, this study further compares the inference efficiency of combined(our CSR and TTS) and unified CSV2S. As an end-to-end CSV2S generation framework, our UniCUE significantly reduces the inference time (4.15s), achieving a 40% reduction compared to the traditional CSR + TTS serial approach (6.93s), thereby confirming the advantages of the unified architecture in efficiency optimization.
5. Conclusion
This work introduces UniCUE, the first unified framework that integrates Cued Speech recognition (CSR) and video-to-speech (CSV2S) generation for Chinese Mandarin CS, addressing the critical challenges of bridging communication barriers between the hearing-impaired and normal-hearing communities. By integrating CS recognition with diffusion-based speech generation, our framework achieves fine-grained visual-semantic alignment and precise temporal synchronization, even with limited CS training data. Key innovations, such as the pose-aware visual processor for detailed visual embedding, the semantic alignment pool for robust visual-text mapping, and the VPA for injecting CSR knowledge into the diffusion decoder, collectively enable speech synthesis that preserves both linguistic accuracy and temporal consistency. Experimental results demonstrate that our unified approach not only outperforms SOTA CSR methods in semantic consistency but also achieves strong performance in temporal alignment. Furthermore, the expanded Chinese CS dataset establishes a solid foundation for future research.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 62471420), GuangDong Basic and Applied Basic Research Foundation (2025A1515012296), and CCF-Tencent Rhino-Bird Open Research Fund.References
- (1)
- Cao et al. (2019) Z. Cao, G. Hidalgo Martinez, T. Simon, S. Wei, and Y. A. Sheikh. 2019. OpenPose: Realtime Multi-Person 2D Pose Estimation using Part Affinity Fields. IEEE Transactions on Pattern Analysis and Machine Intelligence (2019).
- Chen et al. (2025a) Liang Chen, Shuai Bai, Wenhao Chai, Weichu Xie, Haozhe Zhao, Leon Vinci, Junyang Lin, and Baobao Chang. 2025a. Multimodal Representation Alignment for Image Generation: Text-Image Interleaved Control Is Easier Than You Think. arXiv:2502.20172 (2025).
- Chen et al. (2025b) Xiaokang Chen, Zhiyu Wu, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, and Chong Ruan. 2025b. Janus-Pro: Unified Multimodal Understanding and Generation with Data and Model Scaling. arXiv preprint arXiv:2501.17811 (2025).
- Chen et al. (2024) Yushen Chen, Zhikang Niu, Ziyang Ma, Keqi Deng, Chunhui Wang, Jian Zhao, Kai Yu, and Xie Chen. 2024. F5-tts: A fairytaler that fakes fluent and faithful speech with flow matching. arXiv:2410.06885 (2024).
- Choromanski et al. (2020) Krzysztof Choromanski, Valerii Likhosherstov, David Dohan, Xingyou Song, Andreea Gane, Tamas Sarlos, Peter Hawkins, Jared Davis, Afroz Mohiuddin, Lukasz Kaiser, et al. 2020. Rethinking attention with performers. arXiv:2009.14794 (2020).
- Chung and Zisserman (2016) J. S. Chung and A. Zisserman. 2016. Out of time: automated lip sync in the wild. In Workshop on Multi-view Lip-reading, ACCV.
- Cornett (1967) R Orin Cornett. 1967. Cued speech. American annals of the deaf (1967), 3–13.
- Dong et al. (2023) Runpei Dong, Chunrui Han, Yuang Peng, Zekun Qi, Zheng Ge, Jinrong Yang, Liang Zhao, Jianjian Sun, Hongyu Zhou, Haoran Wei, et al. 2023. Dreamllm: Synergistic multimodal comprehension and creation. arXiv:2309.11499 (2023).
- Fang et al. (2024) Rongyao Fang, Chengqi Duan, Kun Wang, Hao Li, Hao Tian, Xingyu Zeng, Rui Zhao, Jifeng Dai, Hongsheng Li, and Xihui Liu. 2024. Puma: Empowering unified mllm with multi-granular visual generation. arXiv:2410.13861 (2024).
- Ge et al. (2024a) Yuying Ge, Yizhuo Li, Yixiao Ge, and Ying Shan. 2024a. Divot: Diffusion Powers Video Tokenizer for Comprehension and Generation. arXiv:2412.04432 (2024).
- Ge et al. (2024b) Yuying Ge, Sijie Zhao, Jinguo Zhu, Yixiao Ge, Kun Yi, Lin Song, Chen Li, Xiaohan Ding, and Ying Shan. 2024b. Seed-x: Multimodal models with unified multi-granularity comprehension and generation. arXiv:2404.14396 (2024).
- Gupta et al. (2024) Akshita Gupta, Tatiana Likhomanenko, Karren Dai Yang, Richard He Bai, Zakaria Aldeneh, and Navdeep Jaitly. 2024. Visatronic: A Multimodal Decoder-Only Model for Speech Synthesis. arXiv:2411.17690 (2024).
- Hershey et al. (2017) Shawn Hershey, Sourish Chaudhuri, Daniel PW Ellis, Jort F Gemmeke, Aren Jansen, R Channing Moore, Manoj Plakal, Devin Platt, Rif A Saurous, Bryan Seybold, et al. 2017. CNN architectures for large-scale audio classification. In 2017 IEEE international conference on acoustics, speech and signal processing (ICASSP). IEEE, 131–135.
- Hua et al. (2022) Weizhe Hua, Zihang Dai, Hanxiao Liu, and Quoc Le. 2022. Transformer quality in linear time. In International conference on machine learning. PMLR, 9099–9117.
- Jin et al. (2023) Yang Jin, Kun Xu, Liwei Chen, Chao Liao, Jianchao Tan, Bin Chen, Chenyi Lei, An Liu, Chengru Song, Xiaoqiang Lei, et al. 2023. Unified language-vision pretraining with dynamic discrete visual tokenization. arXiv:2309.04669 (2023).
- Kefalas et al. (2024) Triantafyllos Kefalas, Yannis Panagakis, and Maja Pantic. 2024. Large-scale unsupervised audio pre-training for video-to-speech synthesis. IEEE/ACM Transactions on Audio, Speech, and Language Processing (2024).
- Kilgour et al. (2018) Kevin Kilgour, Mauricio Zuluaga, Dominik Roblek, and Matthew Sharifi. 2018. Fr’echet audio distance: A metric for evaluating music enhancement algorithms. arXiv:1812.08466 (2018).
- Kim et al. (2023) Minsu Kim, Joanna Hong, and Yong Man Ro. 2023. Lip-to-speech synthesis in the wild with multi-task learning. In ICASSP 2023-2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 1–5.
- Lei et al. (2024a) Songju Lei, Xize Cheng, Mengjiao Lyu, Jianqiao Hu, Jintao Tan, Runlin Liu, Lingyu Xiong, Tao Jin, Xiandong Li, and Zhou Zhao. 2024a. Uni-Dubbing: Zero-Shot Speech Synthesis from Visual Articulation. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 10082–10099.
- Lei et al. (2024b) Wentao Lei, Li Liu, and Jun Wang. 2024b. Bridge to Non-Barrier Communication: Gloss-Prompted Fine-grained Cued Speech Gesture Generation with Diffusion Model. arXiv:2404.19277 (2024).
- Leybaert et al. (2010) Jacqueline Leybaert, Mario Aparicio, and Jésus Alegria. 2010. 19 The Role of Cued Speech in Language Development of Deaf Children. The Oxford Handbook of Deaf Studies, Language, and Education, Volume 1 (2010), 276.
- Leybaert and LaSasso (2010) Jacqueline Leybaert and Carol J LaSasso. 2010. Cued speech for enhancing speech perception and first language development of children with cochlear implants. Trends in amplification 14, 2 (2010), 96–112.
- Li et al. (2023) Junnan Li, Dongxu Li, Silvio Savarese, and Steven Hoi. 2023. Blip-2: Bootstrapping language-image pre-training with frozen image encoders and large language models. In International conference on machine learning. PMLR, 19730–19742.
- Li et al. (2024) Yanwei Li, Yuechen Zhang, Chengyao Wang, Zhisheng Zhong, Yixin Chen, Ruihang Chu, Shaoteng Liu, and Jiaya Jia. 2024. Mini-gemini: Mining the potential of multi-modality vision language models. arXiv:2403.18814 (2024).
- Liu et al. (2023) Haohe Liu, Zehua Chen, Yi Yuan, Xinhao Mei, Xubo Liu, Danilo Mandic, Wenwu Wang, and Mark D Plumbley. 2023. Audioldm: Text-to-audio generation with latent diffusion models. arXiv:2301.12503 (2023).
- Liu and Feng (2019) Li Liu and Gang Feng. 2019. A pilot study on mandarin chinese cued speech. American Annals of the Deaf 164, 4 (2019), 496–518.
- Liu et al. (2020) Li Liu, Gang Feng, Denis Beautemps, and Xiao-Ping Zhang. 2020. Re-synchronization using the hand preceding model for multi-modal fusion in automatic continuous cued speech recognition. IEEE Transactions on Multimedia 23 (2020), 292–305.
- Liu et al. (2018) Li Liu, Thomas Hueber, Gang Feng, and Denis Beautemps. 2018. Visual Recognition of Continuous Cued Speech Using a Tandem CNN-HMM Approach.. In Interspeech. 2643–2647.
- Liu et al. (2019) Li Liu, Jianze Li, Gang Feng, and Xiao-Ping Steven Zhang. 2019. Automatic Detection of the Temporal Segmentation of Hand Movements in British English Cued Speech.. In Interspeech. 2285–2289.
- Liu and Liu (2023) Lei Liu and Li Liu. 2023. Cross-modal mutual learning for cued speech recognition. In ICASSP 2023-2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 1–5.
- Liu et al. (2024) Lei Liu, Li Liu, and Haizhou Li. 2024. Computation and parameter efficient multi-modal fusion transformer for cued speech recognition. IEEE/ACM Transactions on Audio, Speech, and Language Processing (2024).
- Liu (2020) Y Liu. 2020. Multilingual denoising pre-training for neural machine translation. arXiv:2001.08210 (2020).
- Maaten and Hinton (2008) Laurens van der Maaten and Geoffrey Hinton. 2008. Visualizing data using t-SNE. Journal of Machine Learning Research (JMLR) 9, 86 (2008), 2579–2605. http://www.jmlr.org/papers/v9/vandermaaten08a.html
- Mi et al. (2025) Zhenxing Mi, Kuan-Chieh Wang, Guocheng Qian, Hanrong Ye, Runtao Liu, Sergey Tulyakov, Kfir Aberman, and Dan Xu. 2025. I Think, Therefore I Diffuse: Enabling Multimodal In-Context Reasoning in Diffusion Models. arXiv:2502.10458 (2025).
- Papadimitriou and Potamianos (2021) Katerina Papadimitriou and Gerasimos Potamianos. 2021. A fully convolutional sequence learning approach for cued speech recognition from videos. In 2020 28th European Signal Processing Conference (EUSIPCO). IEEE, 326–330.
- Podell et al. (2023) Dustin Podell, Zion English, Kyle Lacey, Andreas Blattmann, Tim Dockhorn, Jonas Müller, Joe Penna, and Robin Rombach. 2023. Sdxl: Improving latent diffusion models for high-resolution image synthesis. arXiv:2307.01952 (2023).
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-resolution image synthesis with latent diffusion models. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 10684–10695.
- Rong and Liu (2024) Yan Rong and Li Liu. 2024. Seeing your speech style: A novel zero-shot identity-disentanglement face-based voice conversion. arXiv preprint arXiv:2409.00700 (2024).
- Sun et al. (2024) Quan Sun, Yufeng Cui, Xiaosong Zhang, Fan Zhang, Qiying Yu, Yueze Wang, Yongming Rao, Jingjing Liu, Tiejun Huang, and Xinlong Wang. 2024. Generative multimodal models are in-context learners. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 14398–14409.
- Sun et al. (2023) Quan Sun, Qiying Yu, Yufeng Cui, Fan Zhang, Xiaosong Zhang, Yueze Wang, Hongcheng Gao, Jingjing Liu, Tiejun Huang, and Xinlong Wang. 2023. Emu: Generative pretraining in multimodality. In The Twelfth International Conference on Learning Representations.
- Wang et al. (2020) Sinong Wang, Belinda Z Li, Madian Khabsa, Han Fang, and Hao Ma. 2020. Linformer: Self-attention with linear complexity. arXiv:2006.04768 (2020).
- Wu et al. (2024a) Chengyue Wu, Xiaokang Chen, Zhiyu Wu, Yiyang Ma, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, Chong Ruan, et al. 2024a. Janus: Decoupling visual encoding for unified multimodal understanding and generation. arXiv:2410.13848 (2024).
- Wu et al. (2023) Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua. 2023. Next-gpt: Any-to-any multimodal llm. arXiv:2309.05519 (2023).
- Wu et al. (2024b) Yecheng Wu, Zhuoyang Zhang, Junyu Chen, Haotian Tang, Dacheng Li, Yunhao Fang, Ligeng Zhu, Enze Xie, Hongxu Yin, Li Yi, et al. 2024b. Vila-u: a unified foundation model integrating visual understanding and generation. arXiv:2409.04429 (2024).
- Yemini et al. (2023) Yochai Yemini, Aviv Shamsian, Lior Bracha, Sharon Gannot, and Ethan Fetaya. 2023. LipVoicer: Generating Speech from Silent Videos Guided by Lip Reading. arXiv:2306.03258 (2023).
- Yu et al. (2023) Lili Yu, Bowen Shi, Ramakanth Pasunuru, Benjamin Muller, Olga Golovneva, Tianlu Wang, Arun Babu, Binh Tang, Brian Karrer, Shelly Sheynin, et al. 2023. Scaling autoregressive multi-modal models: Pretraining and instruction tuning. arXiv:2309.02591 2, 3 (2023).
- Zhang et al. (2023) Yuxuan Zhang, Lei Liu, and Li Liu. 2023. Cuing without sharing: A federated cued speech recognition framework via mutual knowledge distillation. In Proceedings of the 31st ACM International Conference on Multimedia. 8781–8789.
- Zhen et al. (2022) Q Zhen, W Sun, H Deng, D Li, Y Wei, B Lv, J Yan, L Kong, and Y Zhong. 2022. cosFormer: Rethinking softmax in attention. In International Conference on Learning Representations.
- Zhou et al. (2024) Chunting Zhou, Lili Yu, Arun Babu, Kushal Tirumala, Michihiro Yasunaga, Leonid Shamis, Jacob Kahn, Xuezhe Ma, Luke Zettlemoyer, and Omer Levy. 2024. Transfusion: Predict the next token and diffuse images with one multi-modal model. arXiv:2408.11039 (2024).

### v3

@@
 Abstract.
-Cued Speech (CS) enhances lipreading through hand coding, providing precise speech perception support for the hearing-impaired.
-CS Video-to-Speech generation (CSV2S) task aims to convert the CS visual expressions (CS videos) of hearing-impaired individuals into comprehensible speech signals.
-Direct generation of speech from CS video (called single CSV2S) yields poor performance due to insufficient CS data.
-Current research mostly focuses on CS Recognition (CSR), which convert video content into linguistic text.
-Based on this, one straightforward way of CSV2S is to combine CSR with a Text-to-Speech system.
-This combined architecture relies on text as an intermediate medium for stepwise cross-modal alignment, which may lead to error propagation and temporal misalignment between speech and video dynamics.
-To address these challenges, we propose a novel approach that directly generates speech from CS videos without relying on intermediate text.
-Building upon this, we propose UniCUE, the first unified framework for CSV2S, whose core innovation lies in the integration of the CSR task that provides fine-grained visual-semantic information to facilitate speech generation from CS videos.
-More precisely, (1) a novel fine-grained semantic alignment pool to ensure precise mapping between visual features and speech contents; (2) a VisioPhonetic adapter to bridge cross-task representations, ensuring seamless compatibility between two distinct tasks (i.e., CSV2S and CSR); (3) a pose-aware visual processor is introduced to enhance fine-grained spatiotemporal correlations between lip and hand movements in CS video.
-Experiments on our new established Chinese CS dataset (14 cuers111Cuer represents the person who perform CS.: 8 hearing-impaired and 6 normal-hearing) show that our UniCUE significantly reduces Word Error Rate by 78.3% and improves lip-speech synchronization by 32% compared to the single CSV2S.
+Cued Speech (CS) enhances lipreading via hand coding, offering visual phonemic cues that support precise speech perception for the hearing-impaired.
+The task of CS Video-to-Speech generation (CSV2S) aims to convert CS videos into intelligible speech signals.
+Most existing research focuses on CS Recognition (CSR), which transcribes video content into text.
+Consequently, a common solution for CSV2S is to integrate CSR with a text-to-speech (TTS) system.
+However, this pipeline relies on text as an intermediate medium, which may lead to error propagation and temporal misalignment between speech and CS video dynamics.
+In contrast, directly generating audio speech from CS video (direct CSV2S) often suffer from the inherent multimodal complexity and the limited availability of CS data.
+To address these challenges, we propose UniCUE, the first unified framework for CSV2S that directly generates speech from CS videos without relying on intermediate text.
+The core innovation of UniCUE lies in integrating a understanding task (CSR) that provides fine-grained CS visual-semantic cues to to guide the speech generation.
+Specifically, UniCUE incorporates a pose-aware visual processor, a semantic alignment pool that enables precise visual–semantic mapping, and a VisioPhonetic adapter to bridge the understanding and generation tasks within a unified architecture.
+To support this framework, we construct UniCUE-HI, a large-scale Mandarin CS dataset containing 11,282 videos from 14 cuers, including both hearing-impaired and normal-hearing individuals.
+Extensive experiments conducted on this dataset demonstrate that UniCUE achieves state-of-the-art (SOTA) performance across multiple evaluation metrics.
 1. Introduction
-Cued Speech (CS) is an advanced visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010).
+Cued Speech (CS) is an visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010).
 CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (Cornett, 1967; Leybaert and LaSasso, 2010; Leybaert et al., 2010).
-In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels, addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019), as illustrated in Figure 1(a).
-CS Video-to-Speech generation (CSV2S) task aims to convert the visual expressions (CS videos) of hearing-impaired individuals into comprehensible speech signals.
-However, directly constructing an end-to-end CSV2S model faces the following challenges: First, the complex semantic correlation between CS video and semantic content requires precise mapping from visual cues (lip reading and hand coding) to speech, while limited CS dataset scale further constrains model capacity.
-Second, fine-grained visual spatiotemporal modeling is crucial to resolve the natural asynchrony where hand gestures initiate earlier than corresponding lip movements i.e., hand-preceding phenomenon (Liu et al., 2020).
-Third, personalized adaptation must address dual variations: (1) inter-cuer differences in CS expression, and (2) acoustic diversity in cuer identity characteristics.
-To the best of our knowledge, no prior work has addressed CSV2S task.
-Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text(Liu et al., 2020, 2024, 2018, 2019), neglecting the critical need for natural speech generation.
-This gap severely limits real-time communication between hearing-impaired and normal-hearing individuals, particularly in educational and social contexts.
-For example, in social gatherings, normal-hearing individuals may need to quickly and accurately understand and respond to questions or information shared by their hearing-impaired friends.
-However, existing CSR methods can only provide text output, which is insufficient to support natural and smooth interactions.
-Moreover, lip-reading video-to-speech synthesis methods like LipVoicer (Yemini et al., 2023) leverage lip motions alone, and they fail to model the complementary hand coding in CS.
-Thus, the primary motivation of this study is to build the first CSV2S system that directly decoding Chinese CS video into speech, as shown in Figure 1(b).
-Based on the existing CSR methods, a straightforward approach to build CSV2S is to combine CSR with a Text-to-Speech (TTS) model, as shown in Figure 2(a).
-However, this combined architecture has fundamental drawbacks.
-Firstly, the intermediate CS text serve as a medium between visual and speech, inheriting recognition errors from the CSR module, leading to inaccurate speech content.
-Secondly, the CSR stage only extracts semantic text but loses fine-grained temporal information from video, resulting in generated speech lacking temporal alignment.
-Recent advances in multimodal generation have demonstrated that integrating semantic reasoning capabilities from vision-language models (VLMs) can significantly enhance complex generation tasks, such as text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a).
-Inspired by this paradigm, we believe that the fine-grained visual understanding inherent in CSR can serve as a semantic compensator for CSV2S generation.
-As illustrated in Figure 2(b), our framework introduces a shared visual processor that bridges CSR and CSV2S tasks.
-This processor operates as a two-way translator: during CSR, it extracts linguistic semantics from spatiotemporal lip-hand patterns; for CSV2S, it maps these semantics to guide speech synthesis while preserving cuer identity.
-The key innovation lies in explicitly modeling the semantic compensation flow, where CSR-derived phoneme-level constraints reduce potential ambiguities in word-level speech generation, overcoming the challenges of fine-grained semantic correlation modeling.
-To address the above challenges, building upon the semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, Unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently, relying solely on raw video embeddings, UniCUE is equipped with a pose-aware visual processor that simultaneously processes both modalities.
-This allows UniCUE to generate a mixed visual representation by combining video embeddings and pose embeddings.
-This mixed setup not only achieves fine-grained spatiotemporal modeling for hand-preceding phenomenon but also enhances robustness to cuer-specific CS expression.
-Secondly, to enhance the semantic alignment between visual cues and spoken content, we introduce a semantic alignment pool.
-This mechanism aligns the video latent space and the pose latent space with linguistic semantics via contrastive learning, enabling the model to capture complex semantic correlations and improve the accuracy of the generated speech.
-Thirdly, we incorporate the visual processor from CSR into the diffusion-based CSV2S model, using an VisioPhonetic Adapter (VPA) to translate fine-grained representation into diffusion-compatible codes.
-Based on the identity correlation between speech and face portrait (Rong and Liu, 2024), this design enables our diffusion decoder to not only effectively integrate the semantic information conveyed by the visual cues but also capture the unique characteristics of the cuer identity in the visual frames, thereby enhancing accurate cuer-specific speech generation.
-To evaluate UniCUE on hearing-impaired individuals, we collect CS data from 8 hearing-impaired cuers.
-Additionally, we extend the Mandarin Chinese CS dataset (Lei et al., 2024b) by incorporating data from 8 hearing-impaired and 6 normal-hearing cuers (individuals who perform CS), resulting in a total of 11,282 CS videos.
-The results on this dataset demonstrate that our method not only achieves accurate linguistic understanding but also preserves the cuer’s characteristics.
+In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels (as illustrated in Figure 1(a)), addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019).
+CS Video-to-Speech generation (CSV2S) task aims to convert CS videos of into comprehensible speech signals.
+However, directly constructing an end-to-end CSV2S model faces several challenges.
+Firstly, this task involves complex multimodal semantic correlations, requiring precise mapping from visual cues (lip movements and hand coding) to acoustic speech, while the limited scale of existing CS datasets further constrains model capacity.
+Secondly, fine-grained spatiotemporal modeling of visual information is essential to resolve the intrinsic asynchrony, i.e., the hand-preceding phenomenon, where hand cues precede corresponding lip movements (Liu et al., 2020a).
+To the best of our knowledge, the CSV2S task has not been explicitly studied in prior literature.
+Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text (Liu et al., 2020a, 2024, 2018, 2019), neglecting the critical need for natural speech generation.
+This limitation significantly impairs real-time communication between hearing-impaired and normal-hearing individuals, especially in educational and social scenarios.
+For instance, in group conversations, normal-hearing participants must quickly comprehend and respond to questions posed by their hearing-impaired peers.
+Textual output from CSR systems is often insufficient for such natural and smooth interactions.
+Additionally, recent lipreading-based video-to-speech models such as LipVoicer (Yemini et al., 2024) rely solely on lip movements, failing to capture the complementary hand-coded information in CS that conveys critical phonemic distinctions.
+These shortcomings underscore the need for a more comprehensive approach.
+Motivated by this, we aim to develop the first Chinese CSV2S system that directly decodes CS videos into intelligible speech, as illustrated in Figure 1(b).
+A straightforward solution, shown in Figure 2(a), is to combine a CSR model with a Text-to-Speech (TTS) system.
+However, this combined pipeline suffers from two key drawbacks.
+Firstly, the intermediate textual representation introduces error propagation, as misrecognitions in the CSR stage lead to incorrect speech output.
+Secondly, the textual intermediate discards fine-grained spatiotemporal cues in the CS video, resulting in synthesized speech that lacks temporal alignment with the visual input.
+To overcome these challenges, we draw inspiration from recent advances in multimodal learning, where semantic reasoning from vision-language models (VLMs) has shown strong promise in tasks like text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a).
+We hypothesize that the multimodal visual understanding inherent in CSR can serve as a semantic bridge to support more accurate and controllable speech generation in CSV2S.
+As depicted in Figure 2(b), we introduce a unified framework that leverages a shared visual processor to bridge CSR (understanding task) and CSV2S (generation task).
+This processor serves as a two-way translator: during CSR, it extracts linguistic semantics from fine-grained lip-hand motion patterns; in CSV2S, it utilizes these semantics to guide speech generation.
+The core innovation of our framework lies in modeling a semantic compensation flow, where phoneme-level supervision from CSR reduces ambiguity in speech synthesis, enabling more faithful and coherent voice generation under complex multimodal conditions.
+Building upon this semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently and rely on raw video embeddings, UniCUE employs a pose-aware processor that fuses video and pose streams into a mixed representation.
+This enables fine-grained spatiotemporal modeling of the hand-preceding phenomenon and improves generalization to cuer-specific expressive styles.
+Secondly, to enhance the alignment between visual and linguistic semantics, we introduce a semantic alignment pool to map the video and pose latent spaces into a shared textual space using contrastive learning.
+This facilitates cross-modal correlation modeling and improves semantic consistency in the generated speech.
+Thirdly, to unify the understanding and generation tasks, we reuse the CSR visual encoder within our diffusion-based CSV2S decoder and introduce a VisioPhonetic Adapter (VPA) that transforms the visual representations into diffusion-compatible codes.
+This design enables the decoder to effectively incorporate fine-grained semantic information derived from multimodal visual inputs
+To evaluate UniCUE on hearing-impaired individuals, we extend the MCCS dataset (Lei et al., 2024b) by adding data from 8 hearing-impaired and 2 normal-hearing cuers111Cuer means the people who perform CS., forming the Unified-HI Corpus with 14 cuers.
+Experimental results on this dataset demonstrate that UniCUE not only produces accurate and intelligible speech, but also maintains temporal synchronization with the CS video.
 The main contributions of this work can be summarized as:
-We propose the first CSV2S generation framework by constructing a new unified multimodal system that integrating CSR capabilities to enhance CSV2S.
+We propose the first CSV2S framework by constructing a unified multimodal system that integrates CSR capabilities to enhance speech generation.
 We propose a pose-aware visual processor and a semantic alignment pool to enhance fine-grained, semantically aligned visual representations, and introduce an VPA module to convert fine-grained semantic information into understandable coding for the speech synthesis model.
-We establish a new Mandarin Chinese CS dataset comprising 8 hearing-impaired and 6 normal-hearing cuers.
-Experimental results demonstrate that our UniCUE framework outperforms the single CSV2S model in terms of speech consistency, accuracy, and time efficiency.
+We construct a new Mandarin Chinese CS dataset comprising both hearing-impaired and normal-hearing cuers.
+Experimental results demonstrate that our UniCUE outperforms the state-of-the-art (SOTA) methods in terms of speech accuracy, consistency, and quality.
 2. Related Work
 2.1. Video-to-Speech Generation
-V2S involves generating natural-sounding speech synchronized with a silent video of a person talking.
-A key challenge in this area is performing complex visual-speech modeling under the limited data.
-To address this challenge, Uni-Dubbing (Lei et al., 2024a) incorporates modality alignment during the pre-training phase on multimodal datasets and then fine-tunes using both multimodal and single-modality audio data.
+V2S aims to synthesize natural speech aligned with silent talking videos, but is challenged by limited data.
+Uni-Dubbing (Lei et al., 2024a) addresses this via modality-aligned pre-training on multimodal data and fine-tuning with both multimodal and audio-only inputs.
 Similarly, Kefalas et al.
-(Kefalas et al., 2024) leverage abundant audio-only datasets to pre-train the speech encoder and decoder, subsequently fine-tuning with paired speech-video data.
-Instead of relying on external data, some works (Kim et al., 2023; Yemini et al., 2023; Gupta et al., 2024) have found that incorporating text transcripts plays a significant role in enhancing V2S generation.
-For example, Kim et al.
-(Kim et al., 2023) propose guiding the model using multimodal supervision, i.e., text and speech, to complement insufficient word representations via multi-task learning.
-In contrast to directly using text transcripts, we innovatively utilize the shared semantic correlations between visual-text and visual-audio in CSR and CSV2S.
+(Kefalas et al., 2024) pre-train on large audio-only corpora before tuning on paired data.
+Some studies (Kim et al., 2023; Yemini et al., 2024; Gupta et al., 2024) incorporate transcripts to enhance generation.
+Kim et al.
+(Kim et al., 2023) use text-speech supervision to improve word-level representation via multi-task learning.
+Existing V2S methods primarily focus on lip reading.
+However, CS conveys phonemic information through both lip and hand movements.
+Ignoring hand cues results in incomplete visual representations and degraded speech synthesis quality, limiting the applicability of these methods to CS.
+Notably, no prior work has addressed the CSV2S task.
 2.2. Cued Speech Recognition
-CS is an augmented lip-reading system enhanced by hand coding, offering significant support to the hearing-impaired community.
-While CSR task focuses on converting CS videos into linguistic texts, leveraging both lips and hands as distinct modalities that provide complementary information (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021).
-Existing CSR methods often process lip and hand features independently, followed by feature fusion for content recognition (Liu et al., 2020, 2024; Liu and Liu, 2023; Zhang et al., 2023).
-Due to the asynchronous nature of lip and hand movements, achieving effective fusion remains a challenging problem.
+CS augments lip reading with hand coding to support the hearing-impaired.
+The CSR task aims to transcribe CS videos into text by leveraging lips and hands as complementary modalities (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021).
+Most CSR methods extract lip and hand features separately and fuse them for recognition (Liu et al., 2020a, 2024; Liu and Liu, 2023; Zhang et al., 2023).
+Due to the asynchronous nature of these modalities, effective fusion remains challenging.
 Liu et al.
-(Liu et al., 2020) introduced a re-synchronization procedure to align hand features with lip features, addressing the temporal mismatch.
-While a transformer-based cross-modal mutual learning framework has been developed to facilitate multimodal interaction (Liu and Liu, 2023; Liu et al., 2024).
+(Liu et al., 2020a) proposed re-synchronization to align hand with lip features, while transformer-based mutual learning (Liu and Liu, 2023; Liu et al., 2024) improves multimodal interaction.
 Zhang et al.
-(Zhang et al., 2023) further explored privacy risks in the training process by applying federated learning.
-In contrast, our approach takes whole video frames as input that models lip and hand features simultaneously, avoiding the complexities of feature fusion.
-Moreover, we propose a pose-aware visual processor to enhance feature representation between the two modalities, thereby improving recognition accuracy.
+(Zhang et al., 2023) addressed privacy concerns via federated learning.
+In contrast, we directly model lip and hand cues from whole frames, avoiding explicit fusion.
+A pose-aware visual processor is introduced further to enhance cross-modal representation and improve performance.
 2.3. Unified Understanding and Generation
-Recent efforts in unifying understanding and generation tasks can be broadly categorized into two paradigms.
-The first paradigm integrates visual-language understanding with external generative models (e.g., diffusion models) to achieve multimodal generation (Wu et al., 2023; Dong et al., 2023; Jin et al., 2023; Li et al., 2024; Sun et al., 2024; Ge et al., 2024b, a).
-For instance, works like (Jin et al., 2023; Li et al., 2024) leverage large language models (LLMs) for semantic comprehension while relying on diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-quality image synthesis.
-The second paradigm focuses on optimizing a LLMs-based foundation models through next-token prediction objectives for both visual understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2024a; Chen et al., 2025b).
-For instance, Transfusion (Zhou et al., 2024) unifies the processes of image understanding and image generation within a transformer.
-This approach not only inherits the strong comprehension ability of language models but also preserves the detailed information of images, achieving better controllable generation results in text-to-image.
-Despite their success, these frameworks predominantly address visual-text scenarios, leveraging understanding for speech generation is unexplored.
-In contrast to prior works, we present the first unified framework that bridges visual understanding and speech generation.
+Recent advances in unifying understanding and generation tasks fall into two main paradigms.
+The first integrates visual-language understanding with external generative models (e.g., diffusion models) for multimodal generation (Wu et al., 2024a; Dong et al., 2024; Jin et al., 2024; Li et al., 2024; Sun et al., 2024; Ge et al., 2024, 2025).
+For example, (Jin et al., 2024; Li et al., 2024) utilize large language models (LLMs) for semantic understanding and diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-fidelity image synthesis.
+The second paradigm trains LLM-based foundation models via next-token prediction for both vision understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2025; Chen et al., 2025b).
+Transfusion (Zhou et al., 2024), for instance, unifies image understanding and generation within a single transformer, enabling controllable text-to-image synthesis by preserving visual details.
+However, existing approaches mainly focus on visual-text settings, leaving visual-to-speech generation underexplored.
+In this work, we introduce the first unified framework that bridges visual understanding and speech generation.
 3. Method
-3.1. UniCUE
-To achieve accurate CSV2S generation, the proposed framework need simultaneously resolve two critical challenges: (1) semantic understanding of the semantic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment.
-Inspired by the auxiliary benefits of unified understanding and generation for multimodal controllable image synthesis (Wu et al., 2023; Dong et al., 2023), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations.
+3.1. Overview of UniCUE
+To achieve accurate CSV2S generation, the proposed method needs to simultaneously address two critical challenges: (1) semantic understanding of the linguistic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment.
+Inspired by the auxiliary benefits of unified understanding and generation for multi-modal controllable image synthesis (Wu et al., 2024a; Dong et al., 2024), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations.
 As illustrated in Figure 3, the framework operates via two pathways.
 CSR: Fine-grained Visual Cues Understanding.
-Given a CS video and its corresponding pose (extracted via OpenPose (Cao et al., 2019)), a pose-aware visual processor extracts the embedding of the visual mixture .
-The visual embedding is first transcribed into linguistic text via a transformer-based text decoder to model a fine-grained spatiotemporal visual representation space.
-The process can be denoted as: , where is the predicted text tokens.
-The learning objective of the CSR path is to predict text tokens as closely as possible to the ground truth text tokens :
-where and are the -th token of the predicted tokens and ground truth, respectively, and is the total number of tokens.
-By transcribing cross-cuer visual cues into linguistic text, CSR path establishes phoneme-level semantic grounding for the visual embedding , which equips with cuer-adaptive, fine-grained semantic understanding capabilities.
+As the recognition pathway, CSR models fine-grained spatiotemporal visual semantics to transcribe CS videos into linguistic sequences.
+Given a CS video and its corresponding pose maps (extracted via OpenPose (Cao et al., 2019)), we first utilize a pose-aware visual processor to extract multi-modal embeddings , which capture lip and hand motion cues.
+And then is fed into a auto-regressive Transformer-based text decoder , which models long-range dependencies and contextual interactions across the sequence to generate the predicted token sequence: , where denotes the predicted token sequence.
+Unlike prior approaches relying on Connectionist Temporal Classification (CTC) loss (Graves et al., 2006), which predict each token independently and thus limit the model’s ability to capture cross-token dependencies and coarticulatory effects, our method employs an auto-regressive decoder supervised by cross-entropy loss.
+This design allows to generate tokens conditioned on previously generated outputs and spatialtemporal visual cues, which is more suited to modeling the asynchronous and dynamic nature of CS.
+To further enhance both token-level precision and sequence-level linguistic consistency, we employ a hybrid training objective: a masked language modeling loss supervises selectively masked ground-truth tokens to enhance contextual understanding; a sequence-level cross-entropy loss enforces supervision over the full sequence to promote accurate transcription.
+The final training objective for CSR is:
+where denotes the ground-truth token sequence.
+This dual-loss strategy enhances token-level accuracy while preserving global sequence semantics, enabling the model to capture subtle visual-linguistic cues and temporal dynamics inherent in CS videos, thus improving recognition performance and supporting speech synthesis.
 CSV2S: Cuer-specific Speech Synthesis.
-Based on the correlation between face portrait and speech (Rong and Liu, 2024), a latent diffusion model (LDM) (Rombach et al., 2022) is trained to synthesize cuer-specific speech conditioned on the learned visual embedding .
-Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes , which are then corrupted into by sequentially injecting Gaussian noise over time steps:
-where represents the noise level at the time step.
-The noised is then input into the LDM (Rombach et al., 2022) to learn the distribution of latent speech codes under the visual condition .
-The objective of the CSV2S generation path can be formulated as:
-where represents the LDM.
-By training LDM, visual information is interacted with latent speech codes, enabling the generated speech to accurately match the dynamic changes in the visual content while preserving the naturalness of personalized speech.
+To directly synthesize intelligible and personalized speech from CS videos, we formulate speech generation as a conditional denoising process within a latent diffusion model (LDM) (Rombach et al., 2022).
+Since both lip shapes and hand cues in CS convey phonemic content, the speech generation is conditioned a refined visual embedding , which is derived by transforming the CSR multimodal feature via a VisioPhonetic adapter (VPA).
+Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes , which are progressively corrupted with Gaussian noise over steps: where denotes the noise level at timestep .
+The noisy latent then denoised by the LDM conditioned on .
+The generation objective is defined as:
+where represents the denoising network. By learning this conditional distribution, our model generates temporally aligned speech that reflects the visual expressions of cuers.
 UniCUE: Unified Understanding and Generation.
-The CSR path learns fine-grained visual embeddings by performing linguistic text recognition, which are then used as conditional inputs for the CSV2S path, enabling CSV2S to leverage the fine-grained visual understanding for accurate speech generation.
-This shared mechanism in our unified framework not only reduces information loss but also avoids error propagation that might arise from intermediate text conversions.
-Therefore, CSV2S can capture cue-specific acoustic features while maintaining linguistic accuracy and temporal consistency, generating cuer-specific speech outputs.
+The CSR pathway learns fine-grained multi-modal visual embeddings through detailed linguistic recognition.
+To bridge the architectural gap between the CSR and the diffusion-based speech generator, we introduce a VPA that transforms into a refined representation .
+These embeddings are subsequently utilized as conditional inputs to the CSV2S pathway, enabling the speech synthesis model to leverage enriched visual understanding for improved generation accuracy.
+By sharing visual feature representations within this unified framework, our approach effectively reduces information loss and mitigates error propagation that often arises from intermediate text conversions.
+As a result, CSV2S is capable of generating cue-specific speech that faithfully preserves linguistic fidelity and temporal alignment, producing personalized and intelligible speech outputs tailored to individual cuers.
 3.2. Pose-aware Visual Processor
-Considering the spatiotemporal correlation between hand coding, lip reading, and their semantic content, both CSV2S and CSR require precise localization of lip and hand keypoints.
-This necessity motivates the visual encoder to capture fine-grained pose features.
-To address this, we develop a pose-aware visual processor to construct mixed visual representations by synergistically integrating pose and video embeddings, as shown in Figure 3.
-Specifically, given the video input and pose input , a tokenizer first splits them into patch tokens.
-Let the ViT-based image encoder be denoted as .
-We extract the video features and pose features from the given video tokens and pose tokens , respectively:
-where and .
-Here, represents the number of visual patch tokens and denotes the dimension of visual embedding.
-The video features and pose features are then concatenated along channel dimension, apply a trainable projector layer to form the mixed visual embedding :
-where MLP refers to the projection layer.
-The CSR path supervises the training of the pose-aware visual processor, guiding it to adapt to cuer-specific CS expressions and learn fine-grained semantic representations.
+Considering the strong spatiotemporal correlation between hand coding, lip movement, and their underlying semantic content, both CSV2S and CSR require accurate modeling of lip and hand motion patterns.
+This necessitates a visual encoder capable of capturing fine-grained and temporally coherent features.
+While video frames offer rich appearance information, they often suffer from redundancy and visual ambiguity.
+In contrast, pose maps provide a compact, structured, and noise-resilient representation of motion dynamics.
+To leverage the complementary strengths of both modalities, we design a pose-aware visual processor that constructs fused visual representations, as shown in Figure 3.
+Specifically, the input to the processor consists of video frames and pose maps , both formatted as tensors of shape , where indicates the frame lengths, and denotes the spatial resolution.
+The processor comprises two main components.
+First, a shared visual encoder extracts spatiotemporal features from both modalities via a sequential architecture: a 2D ResNet backbone extracts frame-wise spatial features, which are stacked along the temporal axis and passed through a 1D temporal convolution to model short-term motion patterns.
+The resulting sequence is then fed into a Transformer encoder to capture long-range temporal dependencies across frames.
+This process yields the video features and pose features , where , with denoting the embedding dimension and being the total number of tokens, where is the number of spatial patches per frame.
+Second, the projection layer integrates the two feature streams.
+The video and pose features are concatenated along the channel dimension and passed through a multi-layer perceptron (MLP), consisting of two linear layers with ReLU activation and LayerNorm, to produce the final mixed visual representation:
+The fused representation serves as a unified visual embedding that drives both recognition and generation pathways.
+In the subsequent modules, this representation is semantically aligned with linguistic content and refined for diffusion-based speech synthesis.
 3.3. Semantic Alignment Pool
-To further enhance fine-grained semantic consistency between visual representation and linguistic content, we propose a semantic alignment mechanism that synchronizes visual, pose, and textual modalities through contrastive learning.
-Specifically, a ViT-based text encoder first projects tokens of ground-truth transcripts into embeddings , sharing the same latent space with visual and pose features extracted by the pose-aware processor.
-We formulate the alignment objective as a multi-modal embedding alignment Loss:
-where Cos(,) denotes the cosine similarity. The total semantic alignment loss is calculated via:
-By individually aligning video and pose features with text embeddings, the model can effectively integrate complementary cues, enhancing its ability to understand visual cues for speech generation.
+To further enhance semantic consistency between visual representation and linguistic content, we introduce a semantic alignment mechanism that aligns video, pose, and textual modalities through contrastive learning.
+Specifically, a ViT-based text encoder encodes the ground-truth transcript tokens into text embeddings .
+The visual features and pose features , extracted by the pose-aware visual processor, are projected into a shared latent space via learnable linear layers.
+The text embedding is similarly projected.
+We adopt a contrastive loss across the batch, treating each video-text and pose-text pair from the same sample as a positive pair, and all others as negatives.
+The loss is denoted as:
+where denotes the cosine similarity between normalized embeddings. The total semantic alignment loss is calculated as:
+By enforcing this high-level alignment, the model is encouraged to extract complementary and discriminative semantics from visual modalities.These aligned features not only enhance linguistic recognition in CSR, but also offer semantically grounded condition for accurate speech synthesis.
 3.4. VisioPhonetic Adapter
-The VisioPhonetic adapter (VPA) acts as a critical translator between the learned pose-aware visual processor in CSR and the diffusion-based speech generator in CSV2S, addressing the representation gap caused by their heterogeneous architectures.
-As illustrated in Figure 4, this lightweight module employs a sequential architecture to progressively refine visual-semantic representations into diffusion-compatible conditioning signal:
+While the CSR-derived embeddings capture rich visual-linguistic semantics, they remain mismatched in format and granularity for direct use in diffusion-based speech generation.
+To bridge this modality gap, we propose the VisioPhonetic Adapter (VPA), which transforms semantically aligned visual features into a phonetic-aware conditioning signal suitable for the LDM.
+As illustrated in Figure 4, this lightweight module employs a sequential architecture to progressively refine visual-semantic representations into a diffusion-compatible conditioning signal:
 which includes two MLPs and a Q-Former-style (Li et al., 2023) cross-attention layer.
-The cross-attention mechanism is specifically designed to bridge the representation gap through learnable semantic queries.
-Formally, given the mixed visual embedding , the process is formulated as: where denotes the learnable query vector initialized from the diffusion model’s latent space, and , , are projection matrices.
-The VPA effectively aligns visual information with the semantic requirements of speech generation, further mapping the semantic connections between hand-lip dynamics, phonetic constraints, and speech.
-temporally related information from the visual features, thereby improving the accuracy, synchronization, and cuer-specific expression in speech.
-3.5. Training and Inference
-The training process is divided into two sequential phases to ensure effective knowledge transfer from visual understanding to speech synthesis.
-Stage 1: CSR focuses on learning fine-grained spatiotemporal representations by jointly optimizing the pose-aware visual processor, the semantic alignment pool, text encoder and decoder.
-This stage trains the model using the cross-entropy loss (as defined in Eq.
-1) and the contrastive alignment loss (as defined in Eq.
-7), formulated as: Stage 2: CSV2S leverages the learned visual representations from Stage 1 to train the VPA network and diffusion model with the objective (denoted in Eq.
-3).
-During inference, the framework operates in a streamlined single-path mode to synthesize speech directly from cued CS videos.
-As illustrated in Figure 5 (b), given an CS video and its corresponding pose, the trained pose-aware visual processor extracts the visual mixture representation , which encodes both semantic content and cuer characteristics.
-A noise sample , sampled from a standard Gaussian distribution , is fed into the trained LDM to predict the speech latent code conditioned on the visual mixture representation over timesteps:
-Subsequently, the predicted speech latent code is decoded into a speech mel-spectrogram via VAE decoder, and a pre-trained Vocoder (Liu et al., 2023) is used to convert the mel-spectrogram into a speech waveform.
-Through this series of steps, the CSV2S system is able to generate natural speech that is consistent with the cuer’s characteristics based on the input CS video and its corresponding pose information.
-| Dataset | Cuers | Sentences | Resolution | FPS |
-| French CS (Liu et al., 2018) | 1 normal-hearing cuer | 238 | 50 | |
-| British CS (Liu et al., 2019) | 1 normal-hearing cuer | 97 | 25 | |
-| Chinese CS (Lei et al., 2024b) | 4 normal-hearing cuers | 4000 | 30 | |
-| Ours (Chinese CS) | 6 normal-hearing and 8 hearing-impaired cuers | 11282 | 30 | |
-| Method | Normal-hearing cuers | Hearing-impaired cuers | |||||
-| WER | LSE-D | LSE-C | FAD | WER | LSE-D | LSE-C | |
-| GT | - | 7.250 | 0.200 | - | - | - | - |
-| CMML (Liu and Liu, 2023)+TTS | 0.614 | 9.327 | 0.124 | 4.908 | 1.102 | 9.825 | 0.090 |
-| CNN+FLASH (Hua et al., 2022)+TTS | 0.956 | 9.773 | 0.090 | 5.747 | 0.998 | 9.943 | 0.070 |
-| CNN+Linformer (Wang et al., 2020)+TTS | 0.991 | 9.874 | 0.082 | 5.944 | 1.092 | 10.014 | 0.067 |
-| CNN+Performer (Choromanski et al., 2020)+TTS | 0.945 | 9.873 | 0.105 | 5.429 | 0.988 | 9.876 | 0.079 |
-| CNN+Cosformer (Zhen et al., 2022)+TTS | 1.069 | 10.063 | 0.076 | 6.355 | 1.229 | 10.115 | 0.064 |
-| EcoCued (Liu et al., 2024)+TTS | 0.672 | 9.495 | 0.114 | 5.326 | 1.621 | 9.846 | 0.087 |
-| Ours (CSR)+TTS | 0.030 | 8.461 | 0.158 | 3.653 | 0.254 | 9.045 | 0.127 |
-| Ours (UniCUE) | 0.056 | 7.432 | 0.194 | 1.749 | 0.293 | 8.716 | 0.153 |
+We use learnable semantic queries , which is initialized by computing the average latent representation from ground-truth mel-spectrograms encoded by the pretrained VAE.
+This provides a phonetic-aware initialization aligned with the diffusion model’s target space.
+These queries act as phonetic slots to extract and reorganize relevant patterns from .
+The cross-attention mechanism operates as: The adapted features serve as the final interface between visual understanding and speech synthesis, ensuring that the generated audio is not only temporally coherent but also linguistically faithful to the video input.
+Dataset | Cuers | Sentences | Character | Word | Resolution | FPS | |
+French CS (Liu et al., 2018) | 1-H | 238 | 12872 | - | 50 | ||
+British CS (Liu et al., 2019) | 1-H | 97 | 2741 | - | 25 | ||
+MCCS (Lei et al., 2024b) | 4-H | 4000 | 131608 | 42256 | 30 | ||
+Unified-HI (Ours) |
+11282 | 350333 | 112664 | 30 | ||
+Method | Normal-hearing cuers | Hearing-impaired cuers | |||||||
+WER | LSE-C | LSE-D | DNSMOS | STOI | WER | LSE-C | LSE-D | DNSMOS | |
+GT | - | 7.274 | 7.314 | 2.79 | - | - | - | - | - |
+CMML | 0.663 | 4.135 | 9.241 | 1.24 | 0.11 | 0.924 | 2.141 | 10.132 | 1.03 |
+EcoCued | 0.657 | 4.327 | 9.146 | 1.28 | 0.12 | 0.917 | 2.165 | 10.079 | 1.07 |
+CSR (Ours) | 0.186 | 4.874 | 9.125 | 2.53 | 0.57 | 0.224 | 3.342 | 9.315 | 2.29 |
+Lip2Speech | 0.803 | 4.215 | 9.367 | 1.03 | 0.05 | 0.989 | 2.424 | 10.816 | 0.02 |
+LipVoicer | 0.754 | 4.361 | 9.226 | 1.12 | 0.08 | 0.971 | 2.623 | 10.517 | 0.04 |
+CSV2S (Ours) | 0.374 | 6.245 | 7.962 | 2.27 | 0.42 | 0.422 | 5.938 | 8.347 | 2.04 |
+UniCUE (Ours) | 0.205 | 6.729 | 7.632 | 2.46 | 0.53 | 0.248 | 6.491 | 8.076 | 2.17 |
 4. Experiment
-4.1. Dataset
-To further enhance the model’s generalization ability and validate its practical effectiveness in facilitating communication for hearing-impaired individuals, we establish a new Chinese CS dataset by incorporating CS data from 2 normal-hearing and 8 hearing-impaired cuers with 4 normal-hearing cuers data in (Lei et al., 2024b), increasing the total number of sentences in the dataset to 11,282.
-A comparison between our dataset and existing CS datasets is presented in Table 1.
-Specifically, we invited 2 normal-hearing cuers and 8 hearing-impaired cuers to simultaneously utter and encode a set of text sentences provided by (Lei et al., 2024b).
-The 10 cuers underwent systematic training to ensure they could perform Mandarin Chinese CS fluently and accurately.
-It is important to note that our dataset was collected with the explicit consent of all cuers and is available for open-source use.
-We use CS data from 6 normal-hearing cuers to train the model, splitting this data into training and test sets in a ratio to ensure effective model training and preliminary validation.
-To more comprehensively evaluate the practical effectiveness in facilitating communication for hearing-impaired individuals, we included all the CS data from 8 hearing-impaired cuers in the test set.
-4.2. Implementation Details
+4.1. Experimental Setting
+Dataset.
+Existing CS datasets are limited to normal-hearing cuers and lack data from hearing-impaired individuals, hindering model generalization to the primary users of assistive communication systems.
+To bridge this gap, we construct a new dataset, the Unified-HI Corpus, which includes CS videos from 8 hearing-impaired and 6 normal-hearing cuers.
+This diverse composition significantly enriches variations in gesture styles, lip movements, and speech patterns.
+The expanded coverage introduces more realistic challenges and better reflects practical use cases, enabling models to capture cue-specific nuances essential for hearing-impaired users.
+A comparison with existing CS datasets is shown in Table 1, and further details on sentence coverage and phoneme distribution are included in Appendix Section 2.
+Due to the noisy speech data from hearing-impaired cuers, we use CS data from 6 normal-hearing cuers for training.
+The data from normal-hearing cuers is split by sentence into training and test sets with a 95:5 ratio to ensure effective training and validation.
+Importantly, all CS data from the 8 hearing-impaired cuers are used in the test set, enabling a robust evaluation of model generalization to this group.
 Architecture Details.
-To construct the CSV2S path, we employ a speech generation pipeline consisting of a VAE encoder and decoder, an LDM, and a Vocoder from AudioLDM (Liu et al., 2023).
-We initialize the pipeline by loading the pre-trained weights of the VAE and LDM.
-For the CSR path, we utilize a conditional translation LLM, MBart (Liu, 2020), which comprises a tokenizer, an image encoder, a text encoder, and a text decoder.
-The pipeline is initialized by loading the pre-trained weights with multi-language corpus222https://huggingface.co/facebook/mbart-large-50.
-Training Details.
-We extract CS pose information—including hands, face, and body—from videos using OpenPose333https://github.com/CMU-Perceptual-Computing-Lab/openpose.
-Raw video frames and pose images are resized to , while the audio is resampled to 16 kHz and converted into mel-spectrograms.
-The entire learning framework is implemented in PyTorch, and all experiments are conducted on an NVIDIA RTX A6000.
-We use the Adam optimizer with a learning rate of for end-to-end training, setting the mini-batch size to 2.
+The CSV2S pathway is entirely built upon the AudioLDM (Liu et al., 2023), including its VAE encoder-decoder, latent diffusion model, and vocoder components.
+For CSR, the Transformer in visual process, tokenizer, text-ViT, and text decoder are initialized from MBart (Liu et al., 2020b).
+Detailed training and inference configurations are provided in Appendix Section 1.
+Evaluation Metrics.
+We evaluate the synthesized speech from three perspectives: linguistic accuracy, temporal synchronization, and speech quality.
+Linguistic accuracy is quantified by the Word Error Rate (WER) between the recognized text and ground truth.
+Temporal synchronization is assessed using SyncNet (Chung and Zisserman, 2016), reporting LSE-D (temporal distance) and LSE-C (confidence score).
+Speech quality is evaluated via STOI (Taal et al., 2010) for intelligibility and DNSMOS (Reddy et al., 2021) for naturalness.
 Comparison Methods.
-To validate the effectiveness of our framework, we conduct extensive comparisons between our method UniCUE (CSV2S with CSR-transferred information) and the following methods: 1) our single CSV2S: an end-to-end trained CSV2S without the shared CSR information; 2) our CSR: comprising of the pose-aware visual processor and text decoder to output CS linguistic text (shown in Figure 3); 3) SOTA CSR methods: including CMML (Liu and Liu, 2023), EcoCued (Liu et al., 2024), and various combinations of CNNs and transformers, such as FLASH (Hua et al., 2022), Linformer (Wang et al., 2020), Performer (Choromanski et al., 2020), and Cosformer (Zhen et al., 2022).
-For fairness, we reproduce these CSR methods on our dataset using the same experimental setup described in their respective papers.
-Since CSR methods only output phoneme-level transcriptions, we propose a two-stage conversion pipeline to synthesis speech, as illustrated in Figure 5 (a).
-The process consists of: 1) converting CSR outputs into Chinese pinyin sequences using rule-based CS patterns, followed by translating the pinyin into Chinese characters with QWen toolkit444https://github.com/QwenLM/Qwen;2) generating the final waveforms using SOTA TTS model F5-TTS (Chen et al., 2024) with reference speech.
-| Method | Normal-hearing cuers | Hearing-impaired cuers | |||||
-| WER | LSE-D | LSE-C | FAD | WER | LSE-D | LSE-C | |
-| GT | - | 7.250 | 0.200 | - | - | - | - |
-| CSR††+TTS | 0.125 | 8.847 | 0.129 | 4.231 | 0.307 | 9.672 | 0.098 |
-| CSR‡+TTS | 0.097 | 8.716 | 0.135 | 4.063 | 0.291 | 9.472 | 0.106 |
-| CSR+TTS | 0.030 | 8.461 | 0.158 | 3.653 | 0.254 | 9.045 | 0.127 |
-| CSV2S† | 0.288 | 8.379 | 0.141 | 3.205 | 0.403 | 8.984 | 9.128 |
-| CSV2S | 0.258 | 8.106 | 0.147 | 2.864 | 0.379 | 8.912 | 0.130 |
-| UniCUE* | 0.104 | 7.741 | 9.172 | 2.219 | 0.332 | 8.884 | 0.139 |
-| UniCUE | 0.056 | 7.432 | 0.194 | 1.749 | 0.293 | 8.716 | 0.153 |
-4.3. Evaluation Metrics
-To evaluate the performance of the synthesized speech, we employ three categories of metrics focusing on linguistic accuracy, audio-visual synchronization, and speech quality.
-For linguistic accuracy, we measure the Word Error Rate (WER) between the recognized text of generated speech and ground truth transcripts.
-The transcripts are normalized by removing punctuation and converting numbers to words prior to alignment.
-To measure the audio-visual synchronization between the generated speech and the corresponding video, we utilize SyncNet (Chung and Zisserman, 2016) and report two key metrics: LSE-D (the temporal distance between audio and video) and LSE-C (the confidence score).
-These values are computed using the pre-trained SyncNet model555https://github.com/joonson/syncnet-python.
-For speech quality, Fréchet Audio Distance (FAD) (Kilgour et al., 2018) is employed to assess the naturalness of generated speech by measuring the distribution divergence between real and synthesized audio in a pretrained VGGish embedding space (Hershey et al., 2017).
+We evaluate our UniCUE against: (1) CSV2S (Ours): direct speech synthesis without CSR assistance; (2) CSR (Ours): including pose-aware visual processor, text encoder and decoder, and semantic alignment pool; (3) CSR methods: CMML (Liu and Liu, 2023) and EcoCued (Liu et al., 2024); (4) V2S methods: Lip2Speech (Choi et al., 2023) and LipVoicer (Yemini et al., 2024).
 4.2. Comparison with SOTA Methods
-We conduct comparison with SOTA methods.
-As shown in Table 2, our framework demonstrates superior performance: 1) Thanks to the pose-aware visual processor and semantic alignment pool, our CSR model achieves WER scores of 0.030 for normal-hearing cuers and 0.254 for hearing-impaired cuers, significantly outperforming previous SOTA CSR methods; 2) Leveraging the strong understanding capabilities of CSR, our UniCUE model also achieves lower WER scores.
-Additionally, our UniCUE surpasses others in FAD, LSE-D, and LSE-C metrics, showcasing high-quality, temporally aligned, and expressive speech generation.
+Quantitative Comparison.
+We compare our framework against SOTA methods, as summarized in Table 2.
+Our CSR model, empowered by the pose-aware visual processor and semantic alignment pool, achieves significantly lower WERs (0.186 for normal-hearing and 0.224 for hearing-impaired cuers), surpassing previous CSR methods.
+Building on this strong semantic understanding, UniCUE outperforms V2S methods across LSE-D, LSE-C, DNSMOS, and STOI metrics, demonstrating superior linguistic accuracy, temporal alignment, and speech quality.
+Qualitative Comparison.
+Mel-spectrogram visualizations (Figure 4 in Appendix) further highlight the advantages of our method, showcasing improved temporal synchronization and clearer acoustic structures compared to others.
+Method | Normal-hearing cuers | Hearing-impaired cuers | |||||||
+WER | LSE-C | LSE-D | DNSMOS | STOI | WER | LSE-C | LSE-D | DNSMOS | |
+GT | - | 7.274 | 7.314 | 2.79 | - | - | - | - | - |
+CSR†† | 0.210 | 4.746 | 9.129 | 2.42 | 0.49 | 0.250 | 3.218 | 9.402 | 2.19 |
+CSR‡ | 0.204 | 4.783 | 9.224 | 2.46 | 0.53 | 0.247 | 3.234 | 9.397 | 2.21 |
+CSR | 0.186 | 4.874 | 9.125 | 2.53 | 0.57 | 0.224 | 3.342 | 9.315 | 2.29 |
+CSV2S† | 0.398 | 6.158 | 8.122 | 2.21 | 0.40 | 0.398 | 5.821 | 8.582 | 1.96 |
+CSV2S | 0.374 | 6.245 | 7.962 | 2.27 | 0.42 | 0.422 | 5.938 | 8.347 | 2.04 |
+UniCUE†† | 0.239 | 6.637 | 7.724 | 2.30 | 0.44 | 0.267 | 6.419 | 8.163 | 2.08 |
+UniCUE‡ | 0.231 | 6.641 | 7.716 | 2.33 | 0.46 | 0.276 | 6.421 | 8.159 | 2.10 |
+UniCUE* | 0.226 | 6.613 | 7.731 | 2.37 | 0.48 | 0.271 | 6.410 | 8.167 | 2.12 |
+UniCUE | 0.205 | 6.729 | 7.632 | 2.46 | 0.53 | 0.248 | 6.491 | 8.076 | 2.17 |
 4.3. Ablation Studies
-To systematically validate the contribution of each component, we conduct controlled ablation studies on both normal-hearing and hearing-impaired test data. The results are listed in Table 3.
-The key findings are analyzed as follows: 1) Unified Training Paradigm.
-Compared with the single CSV2S baseline, our UniCUE reduces WER by 62% (0.056 vs.
-0.258) for normal-hearing cuers and 22.7% (0.293 vs.
-0.379) for hearing-impaired groups.
-This verifies that transferring fine-grained semantic knowledge from CSR to CSV2S through shared visual representation learning effectively alleviates the challenge of complex semantic correlation.
-2) Visual Processor Design.
-We analyze the impact of pose-aware visual processor by comparing CSR‡+TTS vs.
-CSR+TTS and CSV2S† vs.
-CSV2S.
-Models relying solely on raw video features exhibit suboptimal performance due to their limited capacity to capture fine-grained motion patterns.
-This suggests that unprocessed video frames contain substantial interference from redundant visual information.
-By integrating pose estimation features as supplementary inputs, the model gains enhanced fine-grained awareness of motions, leading to significant improvements in both CSR and CSV2S tasks.
-Furthermore, the pose features enhances model robustness against inter-cuer variability.
-As illustrated in Figure 6, our model maintains consistent performance across six normal-hearing cuers when speaking same linguistic content, demonstrating the importance of introducing pose-aware information.
-3) Semantic Alignment Mechanism.
-Disabling the semantic alignment pool (CSR††+TTS) weakens the pose-text correlation, causing 17.9% relative WER increase (0.030 vs.
-0.125).
-The contrastive alignment loss enforces spatiotemporal consistency between visual cues and CS phonemes, which is particularly beneficial for accurate semantic content.
-4) VPA Mechanism.
-Removing the VPA (UniCUE*) results in an 85.7% degradation in FAD (1.749 vs.
-2.219), indicating that the cross-attention-based VPA effectively bridges the representation gap by adaptively selecting critical visual-semantic cues relevant to speech characteristics through learnable query interaction.
-5) Qualitative Analysis.
-To more comprehensively demonstrate the advantages of our unified framework UniCUE, we also provide qualitative comparison of CSR+TTS, CSV2S, and UniCUE.
-As shown in Figure 7, UniCUE achieves more accurate temporal alignment with the ground truth, indicating its superior ability to capture temporal dynamics within the visual modality.
-In contrast, CSR and CSV2S exhibit noticeable temporal drift.
+To verify the contribution of each component, we conduct ablation studies on both normal-hearing and hearing-impaired test data. Results are summarized in Table 3.
+Unified Training Paradigm.
+Compared to direct CSV2S, UniCUE reduces WER by 45% (0.205 vs.
+0.374) on normal-hearing cuers and 41% (0.248 vs.
+0.422) on hearing-impaired cuers.
+These results highlight the benefit of leveraging fine-grained visual semantics from CSR to enhance CSV2S, alleviating the challenge of modeling complex multimodal correlations.
+Visual Processor Design.
+Models that rely solely on raw video features struggle to capture fine-grained motion due to redundant and noisy visual information, resulting in suboptimal performance.
+By incorporating pose cues, our visual processor effectively captures cuer-specific dynamics, leading to significantly improved accuracy and robustness across diverse cuers.
+Semantic Alignment Mechanism.
+Disabling the Semantic Alignment Pool (SAP) degrades visual-semantic consistency, resulting in higher WERs for both CSR and UniCUE.
+This underscores the importance of the alignment in enforcing spatiotemporal coherence between visual cues and phonemic representations for accurate semantic modeling.
+The effectiveness of SAP is further validated by the t-SNE visualizations (Figure 5 in Appendix).
+VisioPhonetic Adapter.
+Removing the VPA results in noticeable degradation in temporal alignment, demonstrating its crucial role in bridging the representation gap between CSR and CSV2S.
+By adaptively selecting and refining fine-grained spatialtemporal visual cues through learnable queries, the VPA enables more accurate and temporally coherent speech synthesis.
+Impact of Hand Cues.
+Removing hand cues leads to substantial performance degradation, particularly for hearing-impaired users who often exhibit limited oral articulation and atypical lip shapes (see Appendix Table 1).
+The results highlight the complementary role of hand gestures in enhancing visual phonemic representations for CS.
+Computational Efficiency. UniCUE achieves faster training convergence and 40% lower inference time than the combined pipeline (see details in Appendix Section 5).
 4.4. User Study
-To holistically evaluate the perceptual quality of synthesized speech, we conduct a user study.
-In this experiment, test samples were randomly selected from the outputs of each cuer, and 10 volunteers were invited to rate the generated speech in three perceptual dimensions on 5-point Likert scales: Accuracy (1: Unintelligible, 5: Perfect), Quality (1: Artificial, 5: Human-like), and Synchronization (1: Desynchronized, 5: Perfectly aligned).
-Four systems were compared: the SOTA CSR baseline (CMML (Liu and Liu, 2023)), our CSR, CSV2S, and the proposed UniCUE.
-Based on the mean and variance values presented in Figure 8, UniCUE demonstrates statistically significant superiority across all metrics.
-These results confirm that the unified framework successfully bridges the gap between visual understanding and speech generation, outperforming both combined pipelines and task-specific baselines in human perception.
-4.7. Discussion and Analysis
-Analysis on Cuer Representation.
-To quantitatively assess the capability in preserving cuer-specific acoustic characteristics, we analyze the latent cuer embeddings extracted from generated speech using Resemblyzer666https://github.com/resemble-ai/Resemblyzer.
-The embeddings are projected into 2D space via t-SNE (Maaten and Hinton, 2008) for visual interpretation.
-As shown in Figure 9, speech samples from the same cuer (e.g., cuer 1 to cuer 6) form tight cluster in the high-dimensional embedding space, while the distances between the cluster centers of different cuers are significantly larger.
-This indicates that the model effectively captures and distinguishes the personalized acoustic features of each cuer.
-Computational Efficiency.
-To quantify the efficiency gains of our unified framework, we conduct time cost evaluations across both training and inference stage.
-(a) Training Efficiency.
-We compared the training performance of CSV2S and UniCUE to evaluate the training efficiency of our proposed unified framework.
-As shown in the Figure 10, the training loss of UniCUE decreases rapidly to a stable value in fewer epochs, indicating that the introduction of fine-grained understanding ability accelerates the learning process of the speech generation model.
-Moreover, the validation loss also decreases concurrently with less fluctuation, demonstrating improved model robustness.
-(b) Real-time Inference.
-To validate the real-time performance of the model, this study further compares the inference efficiency of combined(our CSR and TTS) and unified CSV2S.
-As an end-to-end CSV2S generation framework, our UniCUE significantly reduces the inference time (4.15s), achieving a 40% reduction compared to the traditional CSR + TTS serial approach (6.93s), thereby confirming the advantages of the unified architecture in efficiency optimization.
+To comprehensively assess the perceptual quality of synthesized speech, we conduct a user study involving 20 randomly selected test samples per cuer.
+Twenty volunteers rate the generated speech on three perceptual dimensions using 5-point Likert scales: Accuracy (1: unintelligible, 5: perfectly intelligible), Quality (1: artificial, 5: human-like), and Synchronization (1: desynchronized, 5: perfectly aligned).
+As shown in Figure 5, UniCUE consistently achieves significantly higher scores across all metrics, demonstrating statistically meaningful improvements.
+These findings validate that our unified framework effectively bridges visual understanding and speech generation, delivering superior performance in human perception compared to both modular pipelines and task-specific baselines.
 5. Conclusion
-This work introduces UniCUE, the first unified framework that integrates Cued Speech recognition (CSR) and video-to-speech (CSV2S) generation for Chinese Mandarin CS, addressing the critical challenges of bridging communication barriers between the hearing-impaired and normal-hearing communities.
-By integrating CS recognition with diffusion-based speech generation, our framework achieves fine-grained visual-semantic alignment and precise temporal synchronization, even with limited CS training data.
-Key innovations, such as the pose-aware visual processor for detailed visual embedding, the semantic alignment pool for robust visual-text mapping, and the VPA for injecting CSR knowledge into the diffusion decoder, collectively enable speech synthesis that preserves both linguistic accuracy and temporal consistency.
-Experimental results demonstrate that our unified approach not only outperforms SOTA CSR methods in semantic consistency but also achieves strong performance in temporal alignment.
-Furthermore, the expanded Chinese CS dataset establishes a solid foundation for future research.
+This work introduces UniCUE, the first unified framework for directly generating speech from CS videos.
+By integrating fine-grained visual understanding with diffusion-based speech synthesis, UniCUE produces intelligible speech with precise temporal alignment.
+Key components including the pose-aware visual processor, semantic alignment pool, and VisioPhonetic Adapter, enable effective knowledge transfer from CS recognition (CSR) to CS video-to-speech generation (CSV2S), enhancing both linguistic accuracy and temporal synchronization.
+Additionally, we introduce the UniCUE-HI corpus, a new CS dataset featuring both normal-hearing and hearing-impaired cuers.
+Extensive experiments on this dataset demonstrate that UniCUE state-of-the-art methods across multiple evaluation metrics.
 Acknowledgements.
@@
 Multimodal Representation Alignment for Image Generation: Text-Image Interleaved Control Is Easier Than You Think.
-arXiv:2502.20172 (2025).
+CoRR (2025).
 - Chen et al.
@@
 Janus-Pro: Unified Multimodal Understanding and Generation with Data and Model Scaling.
-arXiv preprint arXiv:2501.17811 (2025).
-- Chen et al.
-(2024) Yushen Chen, Zhikang Niu, Ziyang Ma, Keqi Deng, Chunhui Wang, Jian Zhao, Kai Yu, and Xie Chen.
-2024.
-F5-tts: A fairytaler that fakes fluent and faithful speech with flow matching.
-arXiv:2410.06885 (2024).
-- Choromanski et al.
-(2020) Krzysztof Choromanski, Valerii Likhosherstov, David Dohan, Xingyou Song, Andreea Gane, Tamas Sarlos, Peter Hawkins, Jared Davis, Afroz Mohiuddin, Lukasz Kaiser, et al.
-2020.
-Rethinking attention with performers.
-arXiv:2009.14794 (2020).
+CoRR (2025).
+- Choi et al.
+(2023) Jeongsoo Choi, Minsu Kim, and Yong Man Ro.
+2023.
+Intelligible Lip-to-Speech Synthesis with Speech Units.
+In Proceedings of the Annual Conference of the International Speech Communication Association, INTERSPEECH, Vol.
+2023.
+4349–4353.
 - Chung and Zisserman (2016) J. S. Chung and A. Zisserman. 2016. Out of time: automated lip sync in the wild. In Workshop on Multi-view Lip-reading, ACCV.
@@
 (2024) Runpei Dong, Chunrui Han, Yuang Peng, Zekun Qi, Zheng Ge, Jinrong Yang, Liang Zhao, Jianjian Sun, Hongyu Zhou, Haoran Wei, et al.
-2023.
+2024.
 DreamLLM: Synergistic Multimodal Comprehension and Creation.
-arXiv:2309.11499 (2023).
+In ICLR.
 - Fang et al.
@@
 PUMA: Empowering Unified MLLM with Multi-granular Visual Generation.
-arXiv:2410.13861 (2024).
-- Ge et al. (2024a) Yuying Ge, Yizhuo Li, Yixiao Ge, and Ying Shan. 2024a. Divot: Diffusion Powers Video Tokenizer for Comprehension and Generation. arXiv:2412.04432 (2024).
+CoRR (2024).
 - Ge et al.
-(2024b) Yuying Ge, Sijie Zhao, Jinguo Zhu, Yixiao Ge, Kun Yi, Lin Song, Chen Li, Xiaohan Ding, and Ying Shan.
-2024b.
+(2025) Yuying Ge, Yizhuo Li, Yixiao Ge, and Ying Shan.
+2025.
+Divot: Diffusion powers video tokenizer for comprehension and generation.
+In Proceedings of the Computer Vision and Pattern Recognition Conference.
+13606–13617.
+- Ge et al.
+(2024) Yuying Ge, Sijie Zhao, Jinguo Zhu, Yixiao Ge, Kun Yi, Lin Song, Chen Li, Xiaohan Ding, and Ying Shan.
+2024.
 SEED-X: Multimodal Models with Unified Multi-granularity Comprehension and Generation.
-arXiv:2404.14396 (2024).
+CoRR (2024).
+- Graves et al.
+(2006) Alex Graves, Santiago Fernández, Faustino Gomez, and Jürgen Schmidhuber.
+2006.
+Connectionist temporal classification: labelling unsegmented sequence data with recurrent neural networks.
+In Proceedings of the 23rd international conference on Machine learning.
+ACM, 369–376.
 - Gupta et al.
@@
 arXiv:2411.17690 (2024).
-- Hershey et al.
-(2017) Shawn Hershey, Sourish Chaudhuri, Daniel PW Ellis, Jort F Gemmeke, Aren Jansen, R Channing Moore, Manoj Plakal, Devin Platt, Rif A Saurous, Bryan Seybold, et al.
-2017.
-CNN architectures for large-scale audio classification.
-In 2017 IEEE international conference on acoustics, speech and signal processing (ICASSP).
-IEEE, 131–135.
-- Hua et al. (2022) Weizhe Hua, Zihang Dai, Hanxiao Liu, and Quoc Le. 2022. Transformer quality in linear time. In International conference on machine learning. PMLR, 9099–9117.
 - Jin et al.
-(2023) Yang Jin, Kun Xu, Liwei Chen, Chao Liao, Jianchao Tan, Bin Chen, Chenyi Lei, An Liu, Chengru Song, Xiaoqiang Lei, et al.
-2023.
-Unified language-vision pretraining with dynamic discrete visual tokenization.
-arXiv:2309.04669 (2023).
+(2024) Yang Jin, Kun Xu, Liwei Chen, Chao Liao, Jianchao Tan, Quzhe Huang, Bin Chen, Chengru Song, Dai Meng, Di Zhang, et al.
+2024.
+Unified Language-Vision Pretraining in LLM with Dynamic Discrete Visual Tokenization.
+In ICLR.
 - Kefalas et al.
@@
 IEEE/ACM Transactions on Audio, Speech, and Language Processing (2024).
-- Kilgour et al.
-(2018) Kevin Kilgour, Mauricio Zuluaga, Dominik Roblek, and Matthew Sharifi.
-2018.
-Fr’echet audio distance: A metric for evaluating music enhancement algorithms.
-arXiv:1812.08466 (2018).
 - Kim et al.
@@
 Bridge to non-barrier communication: gloss-prompted fine-grained cued speech gesture generation with diffusion model.
-arXiv:2404.19277 (2024).
+In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence.
+6333–6341.
 - Leybaert et al.
@@
 Mini-Gemini: Mining the Potential of Multi-modality Vision Language Models.
-arXiv:2403.18814 (2024).
+CoRR (2024).
 - Liu et al.
@@
 AudioLDM: Text-to-Audio Generation with Latent Diffusion Models.
-arXiv:2301.12503 (2023).
+In International Conference on Machine Learning.
+PMLR, 21450–21474.
 - Liu and Feng (2019) Li Liu and Gang Feng. 2019. A pilot study on mandarin chinese cued speech. American Annals of the Deaf 164, 4 (2019), 496–518.
 - Liu et al.
-(2020) Li Liu, Gang Feng, Denis Beautemps, and Xiao-Ping Zhang.
-2020.
+(2020a) Li Liu, Gang Feng, Denis Beautemps, and Xiao-Ping Zhang.
+2020a.
 Re-synchronization using the hand preceding model for multi-modal fusion in automatic continuous cued speech recognition.
@@
 IEEE/ACM Transactions on Audio, Speech, and Language Processing (2024).
-- Liu (2020) Y Liu. 2020. Multilingual denoising pre-training for neural machine translation. arXiv:2001.08210 (2020).
-- Maaten and Hinton (2008) Laurens van der Maaten and Geoffrey Hinton.
-2008.
-Visualizing data using t-SNE.
-Journal of Machine Learning Research (JMLR) 9, 86 (2008), 2579–2605.
-http://www.jmlr.org/papers/v9/vandermaaten08a.html
+- Liu et al.
+(2020b) Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan Ghazvininejad, Mike Lewis, and Luke Zettlemoyer.
+2020b.
+Multilingual denoising pre-training for neural machine translation.
+Transactions of the Association for Computational Linguistics 8 (2020), 726–742.
 - Mi et al.
@@
 arXiv:2307.01952 (2023).
+- Reddy et al.
+(2021) Chandan KA Reddy, Vishak Gopal, and Ross Cutler.
+2021.
+DNSMOS: A non-intrusive perceptual objective speech quality metric to evaluate noise suppressors.
+In ICASSP 2021-2021 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP).
+IEEE, 6493–6497.
 - Rombach et al.
@@
 10684–10695.
-- Rong and Liu (2024) Yan Rong and Li Liu. 2024. Seeing your speech style: A novel zero-shot identity-disentanglement face-based voice conversion. arXiv preprint arXiv:2409.00700 (2024).
 - Sun et al.
@@
 In The Twelfth International Conference on Learning Representations.
-- Wang et al. (2020) Sinong Wang, Belinda Z Li, Madian Khabsa, Han Fang, and Hao Ma. 2020. Linformer: Self-attention with linear complexity. arXiv:2006.04768 (2020).
+- Taal et al.
+(2010) Cees H Taal, Richard C Hendriks, Richard Heusdens, and Jesper Jensen.
+2010.
+A short-time objective intelligibility measure for time-frequency weighted noisy speech.
+In 2010 IEEE international conference on acoustics, speech and signal processing.
+IEEE, 4214–4217.
 - Wu et al.
-(2024a) Chengyue Wu, Xiaokang Chen, Zhiyu Wu, Yiyang Ma, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, Chong Ruan, et al.
-2024a.
+(2025) Chengyue Wu, Xiaokang Chen, Zhiyu Wu, Yiyang Ma, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, Chong Ruan, et al.
+2025.
 Janus: Decoupling visual encoding for unified multimodal understanding and generation.
-arXiv:2410.13848 (2024).
-- Wu et al. (2023) Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua. 2023. Next-gpt: Any-to-any multimodal llm. arXiv:2309.05519 (2023).
+In Proceedings of the Computer Vision and Pattern Recognition Conference.
+12966–12977.
+- Wu et al. (2024a) Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua. 2024a. NExT-GPT: Any-to-Any Multimodal LLM. In International Conference on Machine Learning. PMLR, 53366–53397.
 - Wu et al.
@@
 arXiv:2409.04429 (2024).
-- Yemini et al. (2023) Yochai Yemini, Aviv Shamsian, Lior Bracha, Sharon Gannot, and Ethan Fetaya. 2023. LipVoicer: Generating Speech from Silent Videos Guided by Lip Reading. arXiv:2306.03258 (2023).
+- Yemini et al.
+(2024) Yochai Yemini, Aviv Shamsian, Lior Bracha, Sharon Gannot, and Ethan Fetaya.
+2024.
+LipVoicer: Generating Speech from Silent Videos Guided by Lip Reading.
+In The Twelfth International Conference on Learning Representations.
 - Yu et al.
@@
 8781–8789.
-- Zhen et al. (2022) Q Zhen, W Sun, H Deng, D Li, Y Wei, B Lv, J Yan, L Kong, and Y Zhong. 2022. cosFormer: Rethinking softmax in attention. In International Conference on Learning Representations.
 - Zhou et al.
@@
 Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model.
-arXiv:2408.11039 (2024).
+In The Thirteenth International Conference on Learning Representations.